### v1

# Referenced internal-line double copy and application to gauge and gravitational beta functions

Abstract
Following the approach of Refs. Tourkine:2012vx ; Geyer:2015jch , the double-copy-like decomposition of exchanged internal states in the world-line limit of one-loop string amplitudes is systematically formulated and generalized to both bosonic and heterotic string theories. As an application, the one-loop beta functions for the gauge and gravitational coupling constants are investigated by analyzing the low-energy field-theory limit of the corresponding three-point one-loop amplitudes in heterotic string theory under a naive $T^{6}$ compactification. Due to supersymmetry, these beta functions vanish trivially. However, by decomposing the scattering integrand according to the different internal loop-exchanged states, the most general model-independent results are obtained.
1 Introduction
The double copy relation reveals a profound connection between the perturbative scattering amplitudes of gauge theories and gravitational theories, expressed as $\text{GR}=\text{YM}^{2}$ Bern:2019prr . Although the tree-level double copy originates from the KLT relations Kawai:1985xq in string theory, recent advances in quantum field theory—such as the 5-loop calculation Bern:2018jmv and the classical solution double copy Monteiro:2014cda —have extended beyond the current reach of perturbative string theory. Nevertheless, string amplitudes continue to serve as an important laboratory for studying scattering amplitudes.
The main focus of research on the perturbative double copy has traditionally been on the correspondence between external states. However, studies about low energy limit of one loop type II string amplitude in Ref.Tourkine:2012vx and similar work on ambitwistor string Geyer:2015jch have opened new avenues for the double copy structure of internal states exchanged at one loop. [^1] Such decomposition is also extended to the two loop amplitude of ambitwistor string.Geyer:2019hnn . Extending these ideas to the bosonic and heterotic string theories forms the central theme of this paper. Furthermore, by adopting the chiral splitting effective formalismDHoker:1988pdl ; DHoker:1989cxq , one-loop integrands of closed string have manifest double copy relation between the left and right movers, which is extensively used in this paper.
On the other hand, the running coupling constant is among the most important predictions of the quantum field theory (QFT). The most general formula of one-loop QCD beta function is given by Gross:1973id ; Politzer:1973fx :
$\beta=-\frac{\,g^{2}\,}{16\pi^{2}}\left(\!\frac{11}{3}C_{v}-\frac{2}{3}n_{f}C_{f}-\frac{1}{6}n_{s}C_{s}\!\right)\!,$ (1)
where $g$ is the gauge coupling constant, $n_{X}$ and $C_{X}$ with $X=v,f,s$ denotes the numbers and the corresponding quadratic Casimir operators of the representations of the relevant vectors (spin-$1$), fermions (spin-$\frac{1}{2}$) and scalars (spin-0) respectively. The non-Abelian nature of QCD gauge theory results in a negative beta function, leading to the celebrated asymptotic freedom of the running gauge coupling constant. This serves as an ideal test case to examine whether the double copy of the reference internal line yields a correct result.
Moreover, given the non-Abelian nature of general relativity (GR), it is intriguing to investigate the loop corrections to the gravitational beta function arising from the graviton (spin-2), dilaton (spin-0, scalar), antisymmetric tensor field (spin-0, axial scalar in 4D), and gravitino (spin-3/2). At the same time, one may expect that gravity also induces corrections to the gauge beta function, proportional to the gravitational constant $\kappa$, where $\kappa^{2}=32\pi G$ and $G$ is the Newtonian gravitational constant. Since the gauge coupling beta function in four dimensions is dimensionless, one would naively expect the following general form:
$\beta_{\text{gauge}}=-\frac{g^{2}}{\,16\pi^{2}\,}\!\left(\!\frac{11}{3}C_{v}\!-\!\frac{2}{3}n_{f}C_{f}\!-\!\frac{1}{6}n_{s}C_{s}\!\right)\!+\!\frac{\,\kappa^{2}\mu^{2}\,}{16\pi^{2}}a\,$, (2)
where $a$ is a coefficient to be determined by one-loop calculations. A nonvanishing coefficient $a$ would imply a power-law running of the gauge coupling constant due to quantum gravitational effects. The possibility of such gravitational power-law corrections in Eq. (2) was first proposed by Robinson and Wilczek Robinson:2005fj , and had resulted in a long time debate Pietrykowski:2006xy ; Toms:2007sk ; Ebert:2007gf ; He:2010mt ; Anber:2010uj ; Ellis:2010rw . This issue is also addressed in the present work with referenced internal line double copy.
Several works in the literature Minahan:1987ha ; DiVecchia:1996uq ; DiVecchia:1996iz derived the gauge beta function from the low-energy limit of string amplitudes. However, these calculations are restricted to specialized string models with fixed field content $\{n_{v},n_{s},n_{f}\}$ . In contrast, the referenced internal-line double-copy approach delivers a significant advantage: it directly yields a universal result valid for arbitrary $\{n_{v},n_{s},n_{f}\}$.
This paper is organized as follow. In Sec.2, the one loop scattering amplitude is systematically decomposed in term of the double copy of internal exchanged particles. The bosonic string is studied in Sec.2.1 and heterotic string is studied in Sec.2.2. We won’t repeat the derivation in Ref.Tourkine:2012vx ; Geyer:2015jch while their main result is included in the decomposition of the supersymmetric right movers of heterotic string. For the bosonic left movers of heterotic string, the fermionic realization is discussed alone. Then the application to beta function calculation is discussed in Sec.3. We present the general structure of three gluons and three graviton one loop amplitude in Sec.3.1 and show how extract Feynman diagrams from different regions of moduli space in Sec.3.2. Later, explicit examples are calculated, including the beta function of gauge coupling constant in Sec.3.3, gravitational beta function in Sec.3.4 and gravitational correction to gauge coupling in Sec.3.5. Finally, the conclusions are summarized in Sec.4.
2 The decomposition of one-loop scattering integrand
2.1 Bosonic string
2.1.1 Structure of one loop amplitude
We firstly try to generalize the decomposition method in Tourkine:2012vx ; Geyer:2015jch to the bosonic closed string amplitude. Because we focus on the field theory limit, only graviton scattering amplitudes are accounted.
The un-integrated and integrated vertex operators of graviton are given by
$V_{i}$ $=\zeta\cdot\partial X(z)\xi\cdot\bar{\partial}X(\bar{z})e^{ik\cdot X(z_{i},\bar{z}_{i})},$ (3)
$U_{i}$ $=\int d^{2}z\zeta\cdot\partial X(z)\xi\cdot\bar{\partial}X(\bar{z})e^{ik\cdot X(z_{i},\bar{z}_{i})},$ (4)
where $\{z,\bar{z}\}$ are the complex coordinates on the world-sheet, $X^{\mu}$ is the spacetime coordinates, $\zeta$ and $\xi$ are polarization vectors.
Then the one loop scattering amplitude can be given by
$\mathcal{M}_{(1)}^{n}=\int_{\mathcal{F}}\frac{d^{2}\tau}{\tau^{d/2}|\eta(\tau)|^{48}}Z_{e}(\tau,\bar{\tau})\langle V_{1}U_{2}...U_{n}\rangle,$ (5)
where without loss of generality, we fix $z_{1}=0$. Here $\mathcal{F}$ denotes the fundamental region of the $\mathbb{SL}(2,Z)$ modular group, $\tau=\tau_{1}+i\tau_{2}$ is the moduli parameter of the one loop world-sheet and $Z_{e}(\tau,\bar{\tau})$ is the corrective factor due to compactification. For instance, we consider string propagates in $R^{1,d-2}\otimes T^{26-d}$ spacetime with radius $R_{i},i=1,...,(26-d)$ and
$Z_{e}(\tau,\bar{\tau})=\prod_{i=1}^{26-d}\sum_{n,w}e^{-\pi\tau_{2}(\frac{\alpha^{\prime}n_{i}^{2}}{R_{i}^{2}}+\frac{w_{i}^{2}R_{i}^{2}}{\alpha^{\prime}})+2\pi inw\tau_{1}}.$ (6)
However, the correlation function $\langle V_{1}U_{2}...U_{n}\rangle$ are not a analytic function in general. Because the scalar propagator on torus is given by
$G(z,\bar{z})=G(z)+\bar{G}(\bar{z})+\frac{2\pi y}{\tau_{2}},$ (7)
where $z=x+iy,x,y\in R$ and we define the chiral scalar propagator $G(z)=-\log\theta_{1}(z,\tau)$. The non-holomorphic part $\frac{2\pi y}{\tau_{2}}$ destroys the double copy structure at one loop level. To restore such double copy structure, we adopt the chiral-splitting effective formalism DHoker:1989cxq ; DHoker:1988pdl and
$\langle V_{1}U_{2}...U_{n}\rangle=\tau_{2}^{d/2}\int d^{d}lf(\zeta_{i},k_{i},l,z,\tau)\bar{f}(\xi_{i},k_{i},l,\bar{z},\bar{\tau})|\mathcal{J}_{n}|^{2}.$ (8)
Here we introduce the loop momentum $l$, the chiral Koba-Nielsen factor $\mathcal{J}_{n}$ is given by
$\mathcal{J}_{n}=$ $\exp[\frac{\pi i\tau\alpha^{\prime}l^{2}}{2}+\pi i\sum_{i}l\cdot k_{i}z_{i}+\sum_{i<j}\frac{\alpha^{\prime}k_{i}\cdot k_{j}}{2}\log\theta_{1}(z_{ij},\tau)],$ (9)
and $f(z)$ ($\bar{f}(\bar{z})$) is a function of $\partial^{1,2}G(z)$ and momentum $k_{i}$ and polarization vectors $\zeta_{i}$ ($\xi_{i}$) of external states. One can see that the double copy structure is restored after introducing the loop momentum $l$ with the cost of non-manifest modular symmetry. More serious loop level KLT relation Stieberger:2023nol was also formulated under this formalism.
Next, we turn to the world-line limit $\tau_{2}\rightarrow\infty$. Because we are interested in the model-independent part, the factor $Z_{e}$ is neglected in the later calculation. It is conventional to expand the above integrand in term of $q=e^{i\pi\tau}$ as well as $\bar{q}=e^{-i\pi\tau}$ and only the coefficient of $q^{n}\bar{q}^{m}$ indicates the mass exchanged states as shown as Table.1.
There is a subtle problem that for given loops in Feynman diagrams or holes on the world sheets, which internal exchanged state corresponds to the $q\bar{q}$ expansion. Recall that in the operator formalismGreen:2012pqa , the one loop open string planar amplitude is proportional to
$A_{n}^{1}\sim\int dp^{d}Tr[\Delta_{1}V_{1}(1)\Delta_{2}V_{2}(1)\Delta_{3}...V_{n}(1)],$ (10)
where $\Delta_{i}=\int_{0}^{1}x_{i}^{L_{0}-2}dx_{i}$ is the world sheet propagator. However, using $x^{L_{0}}V(1)=V(x)x^{L_{0}}$, above equation becomes
$A_{n}^{1}\sim\int dp^{d}\prod_{i}\int_{0}^{1}dx_{i}Tr[V_{1}(x_{1})V_{2}(x_{1}x_{2})...V_{n}(x_{1}x_{2}...x_{m})w^{L_{0}-2}],w=x_{1}x_{2}...x_{m}.$ (11)
In the way, a special referenced propagator is chosen to simplify the detailed calculation. What more, the partition arises form the chosen propagator after expanding the world sheet fields into different modes and performing the loop integral $\int dp^{d}$. Consequently, the propagator corresponds to partition function is the one near the vertex operator fixed to $z=0$ or $\bar{z}=0$, which is referred as referenced internal line in this paper.
Even the physic meaning is much more clear in operator formalism, the path integral formalism is better to calculation for the benefit of conformal symmetry. We will not dive into the detail of this traditional method. In addition, we can tackle the holomorphic part and the anti-holomorphic part separately with the chiral splitting effective (path integral) formalism.
$q^{-2}\bar{q}^{-2}$ tachyon
$q^{0}\bar{q}^{0}$ massless states
$q^{2n}\bar{q}^{2m},m=n>1$ Regge states
Table 1: Unphysical states with $n\neq m$ are removed by the level matching condition, which is realized by the real part integral of $\tau_{1}$.
For later convenience, we define the chiral integrand
$W(z)=f(\zeta_{i},k_{i},l,\tau)\mathcal{J}_{n}=\sum_{n}W_{n}(z)q^{n},$ (12)
which can be analytic to the planar diagram integrand of open string scattering amplitude. By the way, We must emphasize that in the original formulation of Ref.Tourkine:2012vx ; Geyer:2015jch , the $\mathcal{J}_{n}$ and $\bar{\mathcal{J}}_{n}$ are absent. While we will see that the necessity to include $\mathcal{J}_{n}$ and $\bar{\mathcal{J}}_{n}$ in bosonic string as well as the 26d left movers of heterotic string.
2.1.2 Decomposition of the scattering integrand
In this section, we focus on the massless states, so $Z_{e}$ can be neglect. Recall that Dekekind eta function has the following property
$\frac{1}{\eta(\tau)^{24}}=\frac{1}{q^{2}}+24+\mathcal{O}(q^{2}),$ (13)
Because the partition function counts physical spectrum, we can split the integrand in terms of exchanged particles,
$I_{L}^{\text{tachyon}}$ $=W_{0}(z),$ (14)
$I_{L}^{\text{massless}}$ $=W_{2}(z)+24W_{0}(z).$ (15)
And because tachyon is scalar, one naively conjecture that
$I_{L}^{\text{scalar}}$ $=n_{s}W_{0}(z),$ (16)
$I_{L}^{\text{vector}}$ $=W_{2}(z)+(d-2)W_{0}(z).$ (17)
Here we generalize the number of scalar from 26d26-d to a general number $n_{s}$ because we are only interested in the field theory limit. However, it turns out that above conjecture only valid for pure massless graviton scattering amplitude.
It is noticeable that any graviton vertex operator can be generated by the multi-linear part of
$\tilde{V}=e^{ik.X+i\zeta_{1}.\partial X+i\xi_{1}.\bar{\partial}X}=e^{i\tilde{k}\cdot X},$ (18)
where we define
$\tilde{k}=k+\zeta_{1}\partial+\xi_{1}\bar{\partial}.$ (19)
Such that
$\tilde{W}(z)$ $=\exp[\frac{\pi i\tau\alpha^{\prime}l^{2}}{2}+\pi i\sum_{i}l\cdot k_{i}z_{i}+\sum_{i<j}\frac{\alpha^{\prime}\tilde{k}_{i}\cdot\tilde{k}_{j}}{2}\log\theta_{1}(z_{ij},\tau)],$
$=\mathcal{J}_{3}\exp[\sum_{i<j}\frac{\alpha^{\prime}\tilde{k}_{i}\cdot\tilde{k}_{j}}{2}\log\theta_{1}(z_{ij},\tau)].$ (20)
Meanwhile $X^{\mu}$ are free fields on the world-sheet, thus we can rewrite $\tilde{W}(z)$ as
$\exp[\sum_{i<j}\frac{\alpha^{\prime}\tilde{k}_{i}\cdot\tilde{k}_{j}}{2}\log\theta_{1}(z_{ij},\tau)]=\prod_{\mu}T_{\mu},$ (21)
where
$T_{\mu}=\exp[\sum_{i<j}\frac{\alpha^{\prime}\tilde{k}_{i}^{\mu}\tilde{k}_{j}^{\mu}}{2}\log\theta_{1}(z_{ij},\tau)]=\sum_{n}T_{\mu}^{(n)}q^{n}.$ (22)
Here $T_{\mu}$ can be identified as the contribution of non-zero modes of $X^{\mu}$. It is straight forward to see that $X^{\mu=1,..,d}$ associates with vector states and $X^{d+1,...,26}$ corresponds to scalar states. In the way,
$I_{L}^{\text{vector}}$ $=\sum_{\mu=1}^{d}(T^{(2)}_{\mu})|_{\text{multi-linear}}+(d-2)W_{0},$ (23)
$I_{L}^{\text{scalar}}$ $=\sum_{\mu=d}^{n_{s}+d}(T^{(2)}_{\mu})|_{\text{multi-linear}}+n_{s}W_{0}.$ (24)
It is easy to see that above formula reduce to our naive conjecture if and only if $\sum_{\mu=d}^{n_{s}+d}\\ (T^{(2)}_{\mu})|_{\text{multi-linear}}=0$. Similar decomposition can be performed for the anti-holomorphic right-moving part.
2.2 Heterotic string
2.2.1 Scattering amplitude with even spin structure
In this subsection, we shall study the scattering amplitude of heterotic string in $R^{1,3}\otimes T^{6}$. In our convention, the right-movers are chosen to be RNS superstringPolchinski:1998rr and the left movers are bosonic string. And the work of Ref.Tourkine:2012vx and Ref.Geyer:2015jch will be revisited when the right-movers are discussed.
The integrated vertex operators of gluon with ghost number 1-1 and 0 are given by
$U^{0,1}_{i}$ $=\int d^{2}zj_{a_{i}}(z)(i\xi_{i}\cdot\bar{\psi})e^{ik_{i}.X},$ (25)
$U^{1,1}_{i}$ $=\int d^{2}zj_{a_{i}}(z)(i\xi_{i}\cdot\partial X+k_{i}\cdot\bar{\psi}\xi_{i}\cdot\bar{\psi})e^{ik.X},$ (26)
where $\psi^{\mu}$ is the world-sheet fermionic fields and $j_{a_{i}}(\bar{z})$ are current operator of WZW modelDiFrancesco:1997nk , which is given by
$j_{a_{i}}=f_{a_{i}bc}\psi_{b}\psi_{c}$ (27)
in the fermionic realization. Similarly, the graviton vertex operator are given by
$U^{0,2}_{i}$ $=\int d^{2}z(i\zeta_{i}\cdot\partial X)(i\xi_{i}\cdot\bar{\psi})e^{ik_{i}\cdot X},$ (28)
$U^{1,2}_{i}$ $=\int d^{2}z(i\zeta_{i}\cdot\partial X)(i\xi_{i}\cdot\bar{\partial}X+k_{i}\cdot\bar{\psi}\xi_{i}\cdot\bar{\psi})e^{ik\cdot X}.$ (29)
For un-integrated vertexes, we continue to use the symbol $V$ respectively.
In any case, the one loop scattering amplitude for even spin structures in the chiral-splitting effective formalismDHoker:1988pdl ; DHoker:1989cxq , are written as
$\mathcal{A}_{n}^{(1)}=$ $\,\frac{~g^{n}_{10}\!}{\,(2\pi)^{3}V\,}\!\!\!\sum_{\alpha_{L},\beta_{L}}\!\sum_{\alpha_{R},\beta_{R}}\!\!\int_{\!\mathcal{F}}\!\!\frac{\,d^{2}\tau\prod_{i}d^{2}z_{i}\,}{(4\pi^{2})^{2}V_{\text{CKV}}}\int d^{4}lP_{L}(\alpha_{L},\beta_{L},\tau)P_{R}(\alpha_{R},\beta_{R},\bar{\tau})$
$\times\!Z_{T^{6}}(\tau,\bar{\tau})W(\alpha_{L},\beta_{L},\tau,l)\bar{W}(\alpha_{R},\beta_{R},\bar{z},\bar{\tau},l)\,$, (30)
where $P_{L}$ and $P_{R}$ are the left- and right-handed partition functions for the given spin structure $(\alpha_{L},\beta_{L})$ and $(\alpha_{R},\beta_{R})$ respectively (excluding compactification), and $Z_{T^{6}}$ is the extra factor induced by $T^{6}$ compactification. $\mathcal{F}$ denotes the fundamental region under modular transformation and is usually chosen as $|\tau|\!\!\geqslant\!1$ and $\frac{1}{2}\!<\!\tau_{1}\!\leqslant\!\frac{1}{2}$, where the moduli parameter $\tau=\tau_{1}+i\tau_{2}$. The quantity $V=(2\pi R)^{6}$ denotes the volume of the extra dimensional space. And $W_{L}$ and $W_{R}$ are the left and right effective correlation under the chiral-splitting formalism with
$W(\alpha_{L},\beta_{L},\tau,l)\bar{W}(\alpha_{R},\beta_{R},\bar{z},\bar{\tau},l)=\langle V_{1}^{1,\star}U_{2}^{1,\star}...U_{n}^{1,\star}\rangle_{[\alpha,\beta]},\star=1,2.$ (31)
To decompose the internal line states, we expand $W^{L}$ and $W^{R}$ in terms of $q$ and $\bar{q}$,
$W(\alpha_{L},\beta_{L},\tau,l)$ $=\sum_{n}W_{n}(\alpha_{L},\beta_{L},\tau,l)q^{n},$ (32)
$\bar{W}(\alpha_{R},\beta_{R},\tau,l)$ $=\sum_{n}\bar{W}_{n}(\alpha_{R},\beta_{R},\bar{\tau},l)\bar{q}^{n}.$ (33)
Now, we can quickly review what people did in Ref.Tourkine:2012vx and Ref.Geyer:2015jch . Recall that the partition function of right-mover can be expanded as
$P_{R}(0,0)$ $=\frac{1}{\bar{q}}+8+\mathcal{O}(\bar{q}),$ (34)
$P_{R}(0,1/2)$ $=-\frac{1}{\bar{q}}+8+\mathcal{O}(\bar{q}),$ (35)
$P_{R}(1/2,0)$ $=-16+\mathcal{O}(\bar{q}).$ (36)
It is observed that the fermions belong to the R sector $(1/2,0)$ and bosons belong to the NS sector (0,0 or 1/2)($0\text${ or }1/2). And the $\frac{1}{\bar{q}}P_{R}(0,0\text{ or }1/2)|_{\bar{q}^{0}}$ should corresponding to exchanging the unphysical tachyon scalar in the loop. So it is argued by Ref. Tourkine:2012vx and Ref.Geyer:2015jch that
$I_{R}^{\text{scalar}}=\frac{n_{s}}{2}(\bar{W}_{0}(0,0)+\bar{W}_{0}(0,1/2)=n_{s}\bar{W}_{0}(0,0).$ (37)
where $n_{s}$ is the number of real scalar particles. In the second equation, we used the fact that the tachyon scalar must be canceled with $\bar{W}_{0}(0,0)=\bar{W}_{n}(0,1/2)$. Even though this result is corrective, we think more reasonable proof should be given by comparing the $\mathbb{Z}_{3}$ orbifold untwisted partition functions  [^2] The generator of $\mathbb{Z}_{3}$ is chosen to be ${\tt g}=e^{i2\pi(\alpha_{1}J_{45}+\alpha_{2}J_{67}+\alpha_{3}J_{89})}e^{i2\pi(\beta_{1}H_{12}+\beta_{2}H_{34}+\beta_{3}H_{56})}.$ (38) where $\{J_{ij}\}$ denote the rotation generators in spacetime and $\{H_{ij}\}$ are the generators of the Cartan subalgebra of $SO(32)$ which act as rotation operators in the internal space.
$\frac{1}{3}\sum_{h}P_{R}^{\tt e\tt h}(0,0)$ $=\frac{1}{\bar{q}}+2+\mathcal{O}(\bar{q}),$ (39)
$\frac{1}{3}\sum_{h}P_{R}^{\tt e\tt h}(0,1/2)$ $=-\frac{1}{\bar{q}}+2+\mathcal{O}(\bar{q}),$ (40)
$\frac{1}{3}\sum_{h}P_{R}^{\tt e\tt h}(1/2,0)$ $=-4+\mathcal{O}(\bar{q}).$ (41)
By counting the degrees of freedom, we can further conclude that
$I_{R}^{\text{vector}}=$ $\frac{1}{2}(\bar{W}_{1}(0,0)-\bar{W}_{1}(0,1/2)$
$\frac{d-2}{2}(\bar{W}_{0}(0,0)+\bar{W}_{0}(0,1/2)$
$=$ $\bar{W}_{1}(0,0)+(d-2)\bar{W}_{0}(0,0),$ (42)
$I_{R}^{\text{fermion}}=$ $-2n_{f}\bar{W}_{0}(1/2,0).$ (43)
where $d$ is the dimensions of flat spacetime and $n_{f}$ is the number of spacetime fermions in 4d. In the toy models studied in this paper, there always is $n_{v}=1$.
It is noteworthy that, aside from $\bar{W}_{1}(0,0)$, all the terms involving $\bar{W}_{0}(0\text{ or }1/2,0)$ in $I_{R}^{\star},\star\in\{\text{vector, scalar and fermion}\}$ are proportional to the degrees of freedom of corresponding states. This observation can be understood as follow. The $\pm\frac{1}{\bar{q}}$ indicates the vacuum in NS sector. And $\bar{W}_{1}(0,0)$ counts exchanging the massless states created by the inserted vertex operators acting on the NS vacuum, while $\bar{W}_{0}(0\text{ or }1/2,0)$ counts exchanging the intrinsic massless modes propagating on the world-sheet. In addition, the Koba-Nielsen factor $\bar{\mathcal{J}}(\bar{z})$ only contributes $\bar{W}_{n}(\alpha_{R},\beta_{R}),n\geq 2$, so that our formulas yield exactly the same result as Ref.Tourkine:2012vx ; Geyer:2015jch where $\bar{\mathcal{J}}(\bar{z})$ is not considered.
The left mover can be identified with bosonic string in the bosonic realization. And the decomposition of internal line is exactly the same as bosonic string as discussed in Sec.2.1. While it is much convenient to use the fermionic realization for gluon and gluino, where the partition function $SO(32)$ heterotic string is given by
$P_{L}(0,0)$ $=\frac{1}{q^{2}}+\frac{32}{q}+504+\mathcal{O}(q),$ (44)
$P_{L}(0,1/2)$ $=\frac{1}{q^{2}}-\frac{32}{q}+504+\mathcal{O}(q),$ (45)
$P_{L}(1/2,0)$ $=65536q^{2}+\mathcal{O}(q^{3}).$ (46)
Because there is no massless state in $(1/2,0)$ sector, we focus on $(0,0)$ and $(0,1/2)$ sector. Thus we can identify
$ce^{ik\cdot X_{L}}\sim$ $\frac{1}{2q^{2}}(W_{0}(0,0)+W_{0}(0,1/2))$ (47)
$c\psi^{a}e^{ik\cdot X_{L}}\sim$ $\frac{32}{2q}(W_{0}(0,0)-W_{0}(0,1/2))$
$+\frac{1}{2q}(W_{0}(0,0)+W_{0}(0,1/2))$ (48)
$c\{\partial X^{\mu},\psi^{a}\psi^{b}\}e^{ik\cdot X_{L}}\sim$ $\frac{32}{2}(W_{1}(0,0)-W_{1}(0,1/2))$
$+\frac{1}{2}(W_{2}(0,0)+W_{2}(0,1/2))$
$+\frac{504}{2}(W_{0}(0,0)+W_{0}(0,1/2))$ (49)
Scalar tachyon $ce^{ik\cdot X_{L}}$ does present in the left-moving spectrum but be removed by the level match condition in heterotic string amplitude. And $SO(32)$ tachyon $c\psi^{a}e^{ik\cdot X_{L}}$ should be canceled by GSO projection, hence we must have
$W_{0}(0,0)$ $=W_{0}(0,1/2),$ (50)
$W_{1}(0,0)$ $=-W_{1}(0,1/2).$ (51)
The main difficulty to extract the gravitational correction is how to split the contribution of $c\partial X^{\mu}e^{ik\cdot X_{L}}$ and $c\psi^{a}\psi^{b}e^{ik\cdot X_{L}}$.
On the other hand, the fermion propagator $S_{\delta}(z,\tau)$ with even spin structure are given by
$S_{\delta}(z,\tau)=\frac{\theta_{\delta}(z,\tau)\theta^{\prime}_{1}(0,\tau)}{\theta_{\delta}(0,\tau)\theta_{1}(z,\tau)},\delta=2,3,4.$ (52)
If $\tau_{2}\rightarrow+\infty$ and $y\sim\log(\tau_{2})$ or $y\simeq\tau_{2}$,
$S_{2}(z,\tau)$ $\simeq\pi\cot(\pi z)-4\pi\sin(2\pi z)q^{2}+\mathcal{O}(q^{3}),$ (53)
$S_{3}(z,\tau)$ $\simeq\pi\csc(\pi z)-4\pi\sin(\pi z)(q-q^{2})+\mathcal{O}(q^{3}),$ (54)
$S_{4}(z,\tau)$ $\simeq\pi\csc(\pi z)+4\pi\sin(\pi z)(q+q^{2})+\mathcal{O}(q^{3}).$ (55)
Besides if $y\sim\frac{1}{\tau_{2}}$, for any spin structure $\delta$,
$S_{\delta}(z,\tau)\sim\frac{1}{z}.$ (56)
One can see that the quadratic Casimir operator arises from the limit $y_{ij}\rightarrow\infty,i\neq j$,
$4\pi^{3}iC_{SO(32)}=32\times 8\pi^{3}i-1\times 16\pi^{3}i+504\times 0.$ (57)
Here we ignore some terms which are proportional to $\frac{1}{w_{21}}$ or $\frac{1}{w_{32}}$, where $w_{ij}=e^{2\pi iz_{ij}}$. Because they can be removed by the real part integral $\int dx_{2}dx_{3}$ after combined with the left movers.
With above evidences, we find that the conventional tropic limit where $q\rightarrow 0$ and $w\rightarrow 0$ only provides the color part of left movers. By the way, the $q^{0}$ term $504=8+496$ counts the massless states in the left moving spectrum, where $8$ counts the $c\partial X^{\mu}e^{ik\cdot X_{L}}$ and $496=\frac{32\times 31}{2}$ counts $c\psi^{a}\psi^{b}e^{ik\cdot X_{L}}$. In the end, we argue that
$I_{L}^{\text{scalar}}=$ $n^{\prime}_{s}W_{0}+\sum_{\mu=d}^{n^{\prime}_{s}+d}(T^{(2)}_{\mu})$, (58)
$I_{L}^{\text{color}}=$ $W_{2}(S)+n_{c}W_{1}+\frac{n_{c}(n_{c}-1)}{2}W_{0},$ (59)
$I_{L}^{\text{vector}}=$ $(d-2)W_{0}+\sum_{\mu=1}^{d}(T^{(2)}_{\mu})$. (60)
Here we define $n_{c}=32$ for $SO(32)$ heterotic string theory and split $W_{2}$ into
$W_{2}=W_{2}(S)+W_{2}(G),$ (61)
where $W_{2}(S)$ and $W_{2}(G)$ denotes the coefficients of $q^{2}$ from the fermion propagator $S_{\alpha_{L},\beta_{L}}(z,\tau)$ and chiral boson propagator $G(z)$ respectively. As the same as bosonic string,
$W_{2}(G)=\sum_{\mu=1}^{d}(T^{(2)}_{\mu})|_{\text{multi-linear}}+\sum_{\mu=d}^{n^{\prime}_{s}+d}(T^{(2)}_{\mu})|_{\text{multi-linear}}.$ (62)
It is noticeable that if all the external states are gluons, $W_{0}$ is always zero for $SO(32)$ heterotic string. This observation guarantees the right moving integrand $W^{R}$ provides the corrective quadratic Casimir operator for the gauge sector.
For consistency test, we can consider the $E8\times E8$ heterotic string and focus on the first $E8$ gauge group,
$P_{L}(0,0)$ $=\frac{1}{q^{2}}+\frac{16}{q}+376+\mathcal{O}(q),$ (63)
$P_{L}(0,1/2)$ $=\frac{1}{q^{2}}-\frac{16}{q}+376+\mathcal{O}(q),$ (64)
$P_{L}(1/2,0)$ $=256+\mathcal{O}(q).$ (65)
It is easy to check that
$4\pi^{3}i\times C_{E8}=8i\pi^{3}\times 16-16i\pi^{3}\times 1+376\times 0-256i\pi^{3}/2.$ (66)
And due to the fact that there is no $q^{-2}$ and $q^{-1}$ terms in $P_{L}(1/2,0)$, $I_{L}^{\text{scalar}}$ and $I_{L}^{\text{vector}}$ are the same as $SO(32)$ heterotic string.
Besides $W_{2}(G)$ is unique and has nontrivial physical meaning. It is interesting to notice that this term is missing in supersymmetrical right movers of heterotic string as well as other superstring theories. So we argue that on one hand, it denotes a correction from $\alpha^{\prime}F^{3}$ effective interaction in the open string planar one-loop amplitude. On the other hand, it describes the quantum correction from gravitational sector to loop scattering processes in heterotic string. Especially, when external states are all gluons or gluinos, this shows how gravity affects non-gravitational perturbative processes. Such loop correction doesn’t present in type II A/B and type I superstring as a simple torus diagram [^3] Branes and open-closed diagrams must be considered.. We will revisit this point at Sec.3.5. This term is also missed in the ambitwistor string descriptionGeyer:2015jch .
2.2.2 Scattering amplitude with odd spin structure
For the odd spin structure, there are a super-moduli (zero mode of 2D gravitino on the worldsheet) and a super conformal killing vector on the torus. So we must insert a picture changing operator and replace one of vertex operators by its partner with ghost number 0,
$\mathcal{A}_{n}^{(1)}(\text{odd})=$ $\,\frac{~g^{n}_{10}\!}{\,(2\pi)^{3}V\,}\!\!\!\!\sum_{\alpha_{R},\beta_{R}}\!\!\int_{\!\mathcal{F}}\!\!\frac{\,d^{2}\tau\prod_{i}d^{2}z_{i}\,}{(4\pi^{2})^{2}V_{\text{CKV}}}\int d^{4}lP_{L}(1/2,1/2,\tau)P_{R}(\alpha_{R},\beta_{R},\bar{\tau})$
$\times\!Z_{T^{6}}(\tau,\bar{\tau})\langle\mathcal{X}V_{1}^{0,\star}U_{2}^{1,\star}...U_{n}^{1,\star}\rangle_{[\alpha,\beta]},\star=1,2\,$, (67)
where $\mathcal{X}$ is the picture changing operator that is given by
$\mathcal{X}=\delta(\beta)S_{F}.$ (68)
The details of these definition can be found at the well known textbookPolchinski:1998rr .
the fermionic propagator on the torus is given by
$S_{1}(z,\tau)=-\partial G(z,\bar{z},\tau)=-\partial\log\theta_{1}(z,\tau)+\frac{2\pi i}{\tau_{2}}y$ (69)
and $\frac{2\pi i}{\tau_{2}}y$ reflects the existence of zero modes of $\psi^{\mu}$ on the torus with odd spin structure. It is known that such CP-violated amplitude reproduce the Chern-Simon term in SYM and SUGRA. Notice that
$P_{L}(1/2,1/2,\tau)=1,$ (70)
which reflects the fact that only massless fermions are exchanged in these diagrams as well as quantum field theory.
At the end of this section, we summary the double copy relation of referenced internal-line of heterotic string in Table.2. It is obvious and trivial to extend this internal loop particle double copy table to other string theories.
$I_{L}^{\text{vector}}$ $I_{L}^{\text{scalar}}$ $I_{L}^{\text{color}}$
$I_{R}^{\text{vector}}$ $h,B,\Phi$} $A^{\prime}$ $A^{a}$
$I_{R}^{\text{scalar}}$ $A^{\prime\prime}$ $\phi$ $\phi^{a}$
$I_{R}^{\text{fermion}}$ $\Psi^{\mu}$ $\psi$ $\lambda^{a}$
Table 2: The double copies of internal loop particles are similar to the conventional double copies of external particles.
3 Application to beta function
3.1 Three-point One-Loop Amplitude of Heterotic String
In this section, we consider the heterotic string under the $T^{6}$ compactification with radii $R_{i},i=4,...,9$. First of all, the three-point scattering amplitude at tree level can be written as:
$\mathcal{A}_{3(0)}$ $=\mathrm{i}\hskip 0.85358pt\hat{g}\hskip 0.85358ptf^{a_{1}a_{2}a_{3}}\hskip 0.85358pt\xi_{1}^{\mu}\xi_{2}^{\nu}\xi_{3}^{\sigma}V_{\mu\nu\sigma}$
$\equiv\mathrm{i}\hskip 0.85358pt\hat{g}\hskip 0.85358ptf^{a_{1}a_{2}a_{3}}\!\left[(\xi_{1}\hskip-0.85358pt\!\cdot\hskip-0.85358pt\xi_{2})(\xi_{3}\!\cdot\hskip-0.85358ptk_{1})\!+\!(\xi_{2}\hskip-0.85358pt\!\cdot\hskip-0.85358pt\xi_{3})(\xi_{1}\!\cdot\hskip-0.85358ptk_{2})\!+\!(\xi_{3}\hskip-0.85358pt\!\cdot\hskip-0.85358pt\xi_{1})(\xi_{2}\!\cdot\hskip-0.85358ptk_{3})\right]\hskip-0.85358pt,$ (71)
where $f^{abc}$ is the structure constant of the non-Abelian gauge group and $V^{\mu\nu\sigma}\!\!=\!\eta^{\mu\nu}k_{1}^{\sigma}\hskip-0.85358pt+\hskip-0.85358pt\eta^{\nu\sigma}k_{2}^{\mu}\hskip-0.85358pt+\hskip-0.85358pt\eta^{\sigma\mu}k_{3}^{\nu}\hskip 0.85358pt$ is the kinematic factor of the cubic gauge interactions. In the above, $\hat{g}\hskip 0.85358pt$ denotes the 10-dimensional Yang-Mills gauge coupling. And the tree-level three-point amplitude of three gravitons can be given by the double copy,
$\mathcal{M}_{3(0)}$ $=\mathrm{i}\hskip 0.85358pt\kappa\hskip 0.85358pt\xi_{1}^{\mu}\xi_{2}^{\nu}\xi_{3}^{\sigma}V_{\mu\nu\sigma}\hskip 0.85358pt\zeta_{1}^{\rho}\zeta_{2}^{m}\zeta_{3}^{n}V_{\rho mn}.$ (72a)
Next, we study the three-point one-loop amplitude. An on-shell three-point scattering amplitude at loop level would vanish identically because the on-shell conditions of massless external states require $k_{i}\!\cdot\hskip-0.85358ptk_{j}\hskip-0.85358pt=\hskip-0.85358pt0\hskip 0.85358pt,i,j=1,2,3$. To compute the physical beta function from the one-loop three-point amplitude, we use the following infrared regularization Minahan:1987ha :
$k_{1}+k_{2}+k_{3}=p\hskip 0.85358pt,$ (73)
with $p^{2}\!=\!0\hskip 0.85358pt$. We may denote $k_{4}\!\equiv\!-p$, giving momentum conservation $k_{1}\!+\!k_{2}\!+\!k_{3}\!+\!k_{4}\!=\!0\hskip 0.85358pt$. And the transversality conditions $\xi_{j}\!\cdot\!k_{j}\!=\zeta_{j}\cdot k_{j}=\!0\hskip 0.85358pt$ hold for $j\!=\!\{1,2,3\}$. Then defining $s_{ij}\!=\!(k_{i}\!+\!k_{j})^{2}$, we reach the following kinematic identity:
$s_{12}+s_{23}+s_{31}=0\,.$ (74)
These constraints preserve conformal invariance of the correlation function. Effectively, this corresponds to a four-point amplitude with one external state in the soft limit. Additionally, we adopt the axial gauge for all the polarization vectors $\xi_{i},i=\{1,2,3\}$, which imposes $\xi_{i}\cdot p=0$.
Then, the one-loop three-gluon scattering amplitude is given by
$\mathcal{A}_{3(1)}=$ $\,\frac{~\hskip 0.85358pt\hat{g}^{3}\!}{\,(2\pi)^{3}V\,}\!\!\!\sum_{\alpha_{L},\beta_{L}}\!\sum_{\alpha_{R},\beta_{R}}\!\!\int_{\!\mathcal{F}}\!\!\frac{\,\mathrm{d}^{2}\tau\hskip 0.85358pt\mathrm{d}^{2}z_{2}\hskip 0.85358pt\mathrm{d}^{2}z_{3}\,}{(4\pi^{2})^{2}}\int d^{4}l|\mathcal{J}_{3}|^{2}$
$\times\!P_{L}(\alpha_{L},\beta_{L},\tau)P_{R}(\alpha_{R},\beta_{R},\bar{\tau})Z_{T^{6}}(\tau,\bar{\tau})\langle V_{1}^{1,1}U_{2}^{1,1}U_{3}^{1,1}\rangle_{[\alpha,\beta]}\,$, (75)
where the correlation function of vertex operators $\langle V_{1}^{1,1}U_{2}^{1,1}U_{3}^{1,1}\rangle_{[\alpha,\beta]}\!$ will be derived in the later Eq.(3.1). Here $W_{L}$ and $W_{R}$ are the left- and right-handed partition functions for the given spin structure [^4] Since there are two periodic directions on a torus, we should specialize two boundary conditions of $\psi^{M}$. Thus, we need to sum over the spin structures $\{\alpha_{R},\beta_{R}\}$ of the right-moving part and $\{\alpha_{L},\beta_{L}\}$ of the left-moving part independently. $(\alpha_{L},\beta_{L})$ and $(\alpha_{R},\beta_{R})$ respectively (excluding compactification), and $Z_{T^{6}}$ is the extra factor induced by $T^{6}$ compactification. $\mathcal{F}$ denotes the fundamental region under modular transformation and is usually chosen as $|\tau|\!\!\geqslant\hskip-0.85358pt\!1$ and $\frac{1}{2}\hskip-0.85358pt\!<\!\hskip-0.85358pt\tau_{1}\hskip-0.85358pt\!\leqslant\hskip-0.85358pt\!\frac{1}{2}\hskip 0.85358pt$, where the moduli parameter $\tau=\tau_{1}+\mathrm{i}\hskip 0.85358pt\tau_{2}$. The quantity $V=(2\pi R)^{6}$ denotes the volume of the extra dimensional space. Finally, $\mathcal{J}_{3}$ is the chiral Koba-Nielsen(KN) factor which is given by
$\mathcal{J}_{3}=$ $\exp[\frac{\pi i\tau\alpha^{\prime}l^{2}}{2}+\pi i\sum_{i}l\cdot k_{i}z_{i}+\sum_{i<j}\frac{\alpha^{\prime}k_{i}\cdot k_{j}}{2}\log\theta_{1}(z_{ij},\tau)]$ (76)
where $l$ is the four dimensional loop momentum. The conventional modular invariant string integrand can be recovered by integrating out $l$.
Given the spin structures $\alpha_{\star},\beta_{\star}]$, we derive the correlation function of the three vertex operators in the chiral splitting effective formalism as follows:
$\langle V_{1}^{1,1}U_{2}^{1,1}U_{3}^{1,1}\rangle_{[\alpha_{\star},\beta_{\star}]}=$ $\frac{\alpha^{\prime 2}}{8}\Big[\Lambda^{1}(\bar{z})+\Lambda^{2}(\alpha_{R},\beta_{R},\bar{z})+\Lambda^{3}(\alpha_{R},\beta_{R},\bar{z})\Big]$
$\Phi_{\alpha_{L}\beta_{L}}\!\!(T^{a_{1}},T^{a_{2}},T^{a_{3}})\hskip 0.85358pt,$ (77)
where $\Lambda^{1}$, $\Lambda^{2}$ and $\Lambda^{3}$ are given by
$\hskip-25.60747pt\Lambda^{1}(\bar{z})=$ $\alpha^{\prime}(\xi_{1}\cdot k_{2}(\bar{\partial}G_{12}-\bar{\partial}G_{13})-2\pi il\cdot\xi_{1})$
$(\xi_{2}\cdot k_{3}(\bar{\partial}G_{12}+\bar{\partial}G_{23})-2\pi il\cdot\xi_{2})$
$\left(\xi_{3}\cdot k_{1}\left(\bar{\partial}G_{23}-\bar{\partial}G_{13}\right)-2\pi il\cdot\xi_{3}\right)$
$-2\xi_{1}\cdot\xi_{2}\bar{\partial}^{2}G_{12}\left(\xi_{3}\cdot k_{1}\left(\bar{\partial}G_{23}-\bar{\partial}G_{13}\right)-2\pi il\cdot\xi_{3}\right)$
$-2\xi_{1}\cdot\xi_{3}\bar{\partial}^{2}G_{13}\left(\xi_{2}\cdot k_{3}\left(\bar{\partial}G_{12}+\bar{\partial}G_{23}\right)-2\pi il\cdot\xi_{2}\right)$
$-2\xi_{2}\cdot\xi_{3}\bar{\partial}^{2}G_{23}\left(\xi_{1}\cdot k_{2}\left(\bar{\partial}G_{12}-\bar{\partial}G_{13}\right)-2\pi il\cdot\xi_{1}\right),$ (78)
$\hskip-25.60747pt\Lambda^{2}(\alpha_{R},\beta_{R},\bar{z})=$ $\alpha^{\prime}\Big[+(k_{1}\cdot k_{2}\xi_{1}\cdot\xi_{2}+\xi_{1}\cdot k_{2}\xi_{2}\cdot k_{3})\xi_{3}\cdot k_{1}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{12}\right)\bar{\partial}G_{13}$
$-(k_{1}\cdot k_{2}\xi_{1}\cdot\xi_{2}+\xi_{1}\cdot k_{2}\xi_{2}\cdot k_{3})\xi_{3}\cdot k_{1}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{12}\right)\bar{\partial}G_{23}$
$-(k_{1}\cdot k_{3}\xi_{1}\cdot\xi_{3}+\xi_{1}\cdot k_{2}\xi_{3}\cdot k_{1})\xi_{2}\cdot k_{3}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{13}\right)\bar{\partial}G_{12}$
$-(k_{2}\cdot k_{3}\xi_{2}\cdot\xi_{3}+\xi_{2}\cdot k_{3}\xi_{3}\cdot k_{1})\xi_{1}\cdot k_{2}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{23}\right)\bar{\partial}G_{12}$
$+(k_{2}\cdot k_{3}\xi_{2}\cdot\xi_{3}+\xi_{2}\cdot k_{3}\xi_{3}\cdot k_{1})\xi_{1}\cdot k_{2}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{23}\right)\bar{\partial}G_{13}$
$-(k_{1}\cdot k_{3}\xi_{1}\cdot\xi_{3}+\xi_{1}\cdot k_{2}\xi_{3}\cdot k_{1})\xi_{2}\cdot k_{3}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{13}\right)\bar{\partial}G_{23}$
$+i2\pi\left(k_{1}\cdot k_{2}\xi_{1}\cdot\xi_{2}+\xi_{1}\cdot k_{2}\xi_{2}\cdot k_{3}\right)l\cdot\xi_{3}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{12}\right)$
$+i2\pi\left(k_{1}\cdot k_{3}\xi_{1}\cdot\xi_{3}+\xi_{1}\cdot k_{2}\xi_{3}\cdot k_{1}\right)l\cdot\xi_{2}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{13}\right)$
$+i2\pi(k_{2}\cdot k_{3}\xi_{2}\cdot\xi_{3}+2\pi\xi_{2}\cdot k_{3}\xi_{3}\cdot k_{1})l\cdot\xi_{1}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{23}\right)\Big],$ (79)
$\Lambda^{3}(\alpha_{R},\beta_{R},\bar{z})$ $=\alpha^{\prime}S_{\alpha_{R}\beta_{R}}\left(\bar{z}_{12}\right)S_{\alpha_{R}\beta_{R}}\left(\bar{z}_{13}\right)S_{\alpha_{R}\beta_{R}}\left(\bar{z}_{23}\right)$
$(+2\alpha^{\prime}\xi_{1}\cdot k_{2}\xi_{2}\cdot k_{3}\xi_{3}\cdot k_{1}+\alpha k_{1}\cdot k_{2}\xi_{1}\cdot\xi_{2}\xi_{3}\cdot k_{1}$
$+\alpha^{\prime}k_{1}\cdot k_{3}\xi_{1}\cdot\xi_{3}\xi_{2}\cdot k_{3}+\alpha k_{2}\cdot k_{3}\xi_{2}\cdot\xi_{3}\xi_{1}\cdot k_{2}).$ (80)
and the quantity $\Phi_{\alpha_{L}\beta_{L}}$ is given by
$\Phi_{\alpha_{L}\beta_{L}}(T^{a_{1}},T^{a_{2}},T^{a_{3}})=S_{\alpha_{L}\beta_{L}}(z_{12})S_{\alpha_{L}\beta_{L}}(z_{31})S_{\alpha_{L}\beta_{L}}(z_{23})Tr[T^{a_{1}}T^{a_{2}}T^{a_{3}}]\hskip 0.85358pt,$ (81)
[…]
And the exact formula of the partition functions are given by
$P_{R}(\alpha_{R},\beta_{R},\bar{\tau})$ $=\frac{1}{2}\rho_{\alpha_{R},\beta_{R}}\frac{\theta^{4}\Big[\begin{aligned} \alpha_{R}\\ \beta_{R}\end{aligned}\Big](\bar{z},\bar{\tau})}{\eta^{12}(\bar{\tau})},$ (82)
$P_{L}(\alpha_{L},\beta_{L},\tau)$ $=\frac{1}{2}\frac{\theta^{16}\Big[\begin{aligned} \alpha_{L}\\ \beta_{L}\end{aligned}\Big](z,\tau)}{\eta^{24}(\tau)},$ (83)
$Z_{T^{6}}(\tau,\bar{\tau})$ $=\prod_{i=1}^{6}\sum_{n,w}e^{-\pi\tau_{2}(\frac{\alpha^{\prime}n_{i}^{2}}{R_{i}^{2}}+\frac{w_{i}^{2}R_{i}^{2}}{\alpha^{\prime}})+2\pi inw\tau_{1}}.$ (84)
Here $\rho_{\alpha_{R},\beta_{R}}$ realizes the GSO projection with $\rho(0,0)=1$ and $\rho(0,1/2)=\rho(1/2,0)=-1$.
However, the spin structure summation (more details are summarized in Sec.A) of the right mover gives that
$\sum_{\alpha_{R},\beta_{R}}P_{R}(\alpha_{R},\beta_{R})S^{n}_{\alpha_{R}\beta_{R}}(z,\tau)$ $=0,n=0,1,2,3,$ (85)
which demonstrates that this one loop three point amplitude vanishes identically, leading to a zero beta function. By counting the degree of freedom, we obtain
$c_{SYM_{N=4}}=c_{A}+4c_{F}+6c_{S}=0\hskip 0.85358pt.$ (86)
This cancellation corresponds to the vanishing of the beta function in $N=4$ SYM, consistent with its conformal invariance.
To explore the gravitational beta function in perturbative theory [^5] Here the stablization of dilation vacuum is not considered., we study the one loop amplitude of three gravitons. The three-graviton one-loop amplitude is given by
$\hskip-8.53581pt\mathcal{M}_{3(1)}\!=$ $\!\!\frac{\hat{\kappa}^{3}}{(2\pi)^{3}V}\!\sum_{\alpha_{L},\beta_{L}}\!\sum_{\alpha_{R},\beta_{R}}\!\!\int_{\mathcal{F}}\!\!\!\frac{\,\mathrm{d}^{2}\tau\hskip 0.85358pt\mathrm{d}^{2}\hskip-0.85358ptz_{2}\hskip 0.85358pt\mathrm{d}^{2}\hskip-0.85358ptz_{3}\,}{(4\pi^{2})^{2}}\int d^{4}l$
$P_{L}(\alpha_{L},\beta_{L},\tau)P_{R}(\alpha_{R},\beta_{R},\bar{\tau})Z_{T^{6}}\langle V_{1}^{1,2}U_{2}^{1,2}U_{3}^{1,2}\rangle_{[\alpha,\beta]}\hskip 0.85358pt.$ (87)
In the way, the torus correlation function $\langle V^{(1,2)}_{1}U^{(1,2)}_{2}U^{(1,2)}_{3}\rangle$ can be given by double copy
$\langle V^{(1,2)}_{1}U^{(1,2)}_{2}U^{(1,2)}_{3}\rangle_{[\alpha,\beta]}=$ $\frac{\alpha^{\prime 4}}{64}\Big[(\Lambda^{1}(\bar{z})+\Lambda^{2}(\alpha_{R},\beta_{R},\bar{z})+\Lambda^{3}(\alpha_{R},\beta_{R},\bar{z}))\Lambda^{b}(z)\Big].$ (88)
Here we define
$\hskip-25.60747pt\Lambda^{b}(z)=$ $\alpha^{\prime}(\zeta_{1}\cdot k_{2}(\partial G_{12}-\partial G_{13})+2\pi il\cdot\zeta_{1})$
$(\zeta_{2}\cdot k_{3}(\partial G_{12}+\partial G_{23})+2\pi il\cdot\zeta_{2})$
$\left(\zeta_{3}\cdot k_{1}\left(\partial G_{23}-\partial G_{13}\right)+2\pi il\cdot\zeta_{3}\right)$
$-2\zeta_{1}\cdot\zeta_{2}\partial^{2}G_{12}\left(\zeta_{3}\cdot k_{1}\left(\partial G_{23}-\partial G_{13}\right)+2\pi il\cdot\zeta_{3}\right)$
$-2\zeta_{1}\cdot\zeta_{3}\partial^{2}G_{13}\left(\zeta_{2}\cdot k_{3}\left(\partial G_{12}+\partial G_{23}\right)+2\pi il\cdot\zeta_{2}\right)$
$-2\zeta_{2}\cdot\zeta_{3}\partial^{2}G_{23}\left(\zeta_{1}\cdot k_{2}\left(\partial G_{12}-\partial G_{13}\right)+2\pi il\cdot\zeta_{1}\right).$ (89)
It is obvious to see that $\Lambda^{b}(z)$ is the complex conjugate of $\Lambda^{1}(\bar{z})$.
3.2 Decomposition of moduli space
The field theory limit is realized by taking the world-sheet torus moduli to its world-line limit $\tau_{2}\!\!\rightarrow\!\!\infty$. It corresponds to shrinking the radius of the closed string or forcing the closed string to propagate up to infinity. After taking the world-line limit $\tau_{2}\!\rightarrow\!\infty$, there are two regions of every $z_{ij}$ which correspond to the Feynman diagrams in the world-line limit, as shown in Fig. 1666For more general string amplitude, more regions should be considered.:
  1. $\bullet$
    $|z_{i}\!-\!z_{j}|\!\sim\!\frac{1}{\,\tau_{2}\,}$, shown as the diagram-(a) in Fig. 1.
  2. $\bullet$
    $|z_{i}\!-\!z_{j}|\!\sim\!\tau_{2}$, shown as the diagram-(b) in Fig.1.
We will refer these diagrams as pinched diagrams and non-pinched diagrams respectively. Different regions of $z_{ij}$ affect how $G(z_{ij},\bar{z}_{ij},\tau)$ and $S_{\alpha,\beta}(z_{ij},\tau)$ behave under the world-line limit $\tau_{2}\rightarrow+\infty$.
[…]
((a))
112233(b)(b)
((b))
Figure 1: One-loop contributions to the three-point amplitude of closed strings. For diagram-(a) the distance between $z_{1}$ and $z_{2}$ are infinitesimally small $\sim\frac{1}{\tau_{2}}$ as $\tau_{2}\!\rightarrow\!\infty$ and form a propagator $\frac{1}{\,s_{12}\,}$, whereas $z_{3}$ is far away from $z_{1}$ and $z_{2}$. There are two similar diagrams with $z_{23}\!\rightarrow\!0$ and $z_{31}\!\rightarrow\!0$ (forming the pole structures $\frac{1}{\,s_{23}\,}$ and $\frac{1}{\,s_{31}\,}$) respectively. For the diagram-(b), all $\{z_{1},z_{2},z_{3}\}$ are far away from each other.
Recall that the bosonic chiral propagator $G(z,\tau)$ is given by
$G(z,\tau)=-\log\theta_{1}(z,\tau).$ (90)
Here we use the conventional abbreviation where $1$ denotes spin structure $(1/2,1/2)$, 22 denotes spin structure $(1/2,0)$, 33 denotes $(0,0)$ and $4$ denotes $(0,1/2)$. In the non-pinched regions,
$G(z)=-\log(\sin(\pi z))-4\sin^{2}(\pi z)q^{2}+\mathcal{O}(q^{4}).$ (91)
Such that $W_{0}$ can be evaluated according to the following world-line limit $\tau_{2}\rightarrow\infty$ and $z\rightarrow i\tau_{2}$,
$G(z,\tau)|_{q^{0}}\simeq$ $i\pi z\Theta(y),$ (92)
$\partial G(z,\tau)|_{q^{0}}\simeq$ $i\pi\Theta(y),$ (93)
$\partial^{2}G(z,\tau)|_{q^{0}}\simeq$ $\pi\delta(y).$ (94)
Here we count $\tau_{2}$ and $z=iy\tau_{2}$ as $q^{0}$. And $W_{2}$ can be evaluated according to
$G(z)|_{q^{2}}$ $=-4\sin^{2}(\pi z)=e^{2i\pi z}+e^{-2i\pi z}-2,$ (95)
$\partial G(z)|_{q^{2}}$ $=-4\pi\sin(2\pi z)=2i(e^{2\pi iz}-e^{-2\pi iz}),$ (96)
$\partial^{2}G(z)|_{q^{2}}$ $=-8\pi^{2}\cos(2\pi z)=-4\pi^{2}(e^{2\pi iz}+e^{-2\pi iz}).$ (97)
Suppose that $y=\Im z>0$, then $e^{2\pi iz}\rightarrow 0$ while $e^{-2\pi iz}\rightarrow\infty$ as $y\rightarrow\tau_{2}$. We can simply ignore $e^{2\pi iz}$ but $e^{-2\pi iz}$ must be handed carefully. It is noticeable that
$e^{-2\pi iz}$ $=e^{2\pi y-2\pi ix},$ (98)
$e^{2\pi i\bar{z}}$ $=e^{2\pi y+2\pi ix}.$ (99)
and we also need to integrate $x$ out. So $e^{-2\pi iz}$ make nonzero contribution if and only if it companies with $e^{2\pi i\bar{z}}$. One can naively embed this factor into a two point amplitude and
$\int_{0}^{\tau_{2}}dye^{4\pi y}e^{-s\alpha^{\prime}|y-y^{2}/\tau_{2}|}\sim\frac{1}{s-\frac{4}{\alpha^{\prime}}}.$ (100)
So $e^{4\pi y}$ corresponds to exchanging a tachyon at other internal line. Because we are interested in the massless exchanged states, this factor $e^{-2\pi iz+2\pi i\bar{z}}=e^{4\pi y}$ can be ignored.
As a consequence, we have the following effect replacement at the non-pinched region,
$G(z)|_{q^{2}}$ $\sim-2,$ (101)
$\partial G(z)|_{q^{2}}$ $\sim 0,$ (102)
$\partial^{2}G(z)|_{q^{2}}$ $\sim 0.$ (103)
On the other hand, at the pinched region $|z|\sim\frac{1}{\tau_{2}}$,the pole in $\log\theta_{1}(z,\tau)$ dominates and
$G(z,\bar{z},\tau)$ $\simeq\log z,$ (104)
$\partial G(z,\tau)$ $\sim\frac{1}{z}.$ (105)
just the same as the scalar propagator on the complex plane. However, this scaling doesn’t imply this amplitude diverges. We change variables of the moduli integral $d^{2}z=dxdy$ to polar coordinates $\rho d\rho d\theta$, where $\rho=|z|$ and $\theta\in[0,2\pi]$. Substituting, we obtain
$\int_{0}^{l}\rho d\rho d\theta\frac{1}{\rho^{2}}\rho^{\alpha^{\prime}X}=\frac{2\pi l^{\alpha^{\prime}X}}{\alpha^{\prime}X}=\frac{2\pi}{\alpha^{\prime}X}.$ (106)
where $X=(\sum_{i\in s}p_{i})^{2}$ for given particles in set $s$ pinching, $l$ is a order one free number and in the second equation, we take the limit $\alpha^{\prime}\rightarrow 0$.
3.3 Gauge beta function
3.3.1 Non-pinched diagram
According to the analysis in Sec.2, the contribution of massless gauge sector to the one loop three gluon amplitude is given by
$\hskip-8.53581pt\mathcal{A}_{3(1)}=\frac{~\hskip 0.85358pt\hat{g}^{3}\!}{\,(2\pi)^{3}V\,}\int_{\mathcal{F}}\!\!\!\frac{\,\mathrm{d}^{2}\tau\hskip 0.85358pt\mathrm{d}^{2}\hskip-0.85358ptz_{2}\hskip 0.85358pt\mathrm{d}^{2}\hskip-0.85358ptz_{3}\,}{(4\pi^{2})^{2}}\int d^{4}l\frac{\alpha^{\prime 2}}{8}(I_{R}^{\text{vector}}+I_{R}^{\text{scalar}}+I_{R}^{\text{fermion}})I_{L}^{\text{color}}.$ (107)
In the non-pinched region, we have two distinguished orders $0=y_{1}<y_{2}<y_{3}<\tau_{2}$ and $0=y_{1}<y_{3}<y_{2}<\tau_{2}$ after gauge fixing $z_{1}=0$. Without loss of generality, we focus on $0=y_{1}<y_{2}<y_{3}$ order in this subsection and $0=y_{1}<y_{3}<y_{2}$ order can be evaluated in the same way.
Thus in the $0=y_{1}<y_{2}<y_{3}$ region,
$\hskip-25.60747pt\Lambda^{1}(\bar{z})=$ $\alpha^{\prime}(\xi_{1}\cdot k_{2}(2\pi i)-2\pi il\cdot\xi_{1})(\xi_{2}\cdot k_{3}(2\pi i)-2\pi il\cdot\xi_{2})$
$\left(\xi_{3}\cdot k_{1}\left(2\pi i\right)-2\pi il\cdot\xi_{3}\right),$ (108)
$\hskip-25.60747pt\Lambda^{2}(\alpha_{R},\beta_{R},\bar{z})=$ $\alpha^{\prime}\Big[-(k_{1}\cdot k_{2}\xi_{1}\cdot\xi_{2}+\xi_{1}\cdot k_{2}\xi_{2}\cdot k_{3})\xi_{3}\cdot k_{1}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{12}\right)\pi i$
$-(k_{1}\cdot k_{2}\xi_{1}\cdot\xi_{2}+\xi_{1}\cdot k_{2}\xi_{2}\cdot k_{3})\xi_{3}\cdot k_{1}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{12}\right)\pi i$
$-(k_{1}\cdot k_{3}\xi_{1}\cdot\xi_{3}+\xi_{1}\cdot k_{2}\xi_{3}\cdot k_{1})\xi_{2}\cdot k_{3}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{13}\right)\pi i$
$-(k_{2}\cdot k_{3}\xi_{2}\cdot\xi_{3}+\xi_{2}\cdot k_{3}\xi_{3}\cdot k_{1})\xi_{1}\cdot k_{2}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{23}\right)\pi i$
$-(k_{2}\cdot k_{3}\xi_{2}\cdot\xi_{3}+\xi_{2}\cdot k_{3}\xi_{3}\cdot k_{1})\xi_{1}\cdot k_{2}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{23}\right)\pi i$
$-(k_{1}\cdot k_{3}\xi_{1}\cdot\xi_{3}+\xi_{1}\cdot k_{2}\xi_{3}\cdot k_{1})\xi_{2}\cdot k_{3}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{13}\right)\pi i$
$+i2\pi\left(k_{1}\cdot k_{2}\xi_{1}\cdot\xi_{2}+\xi_{1}\cdot k_{2}\xi_{2}\cdot k_{3}\right)l\cdot\xi_{3}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{12}\right)$
$+i2\pi\left(k_{1}\cdot k_{3}\xi_{1}\cdot\xi_{3}+\xi_{1}\cdot k_{2}\xi_{3}\cdot k_{1}\right)l\cdot\xi_{2}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{13}\right)$
$+i2\pi(k_{2}\cdot k_{3}\xi_{2}\cdot\xi_{3}+2\pi\xi_{2}\cdot k_{3}\xi_{3}\cdot k_{1})l\cdot\xi_{1}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{23}\right)\Big],$ (109)
$\Lambda^{3}(\alpha_{R},\beta_{R},\bar{z})$ $=\alpha^{\prime}S_{\alpha_{R}\beta_{R}}\left(\bar{z}_{12}\right)S_{\alpha_{R}\beta_{R}}\left(\bar{z}_{13}\right)S_{\alpha_{R}\beta_{R}}\left(\bar{z}_{23}\right)$
$(+2\alpha^{\prime}\xi_{1}\cdot k_{2}\xi_{2}\cdot k_{3}\xi_{3}\cdot k_{1}+\alpha k_{1}\cdot k_{2}\xi_{1}\cdot\xi_{2}\xi_{3}\cdot k_{1}$
$+\alpha^{\prime}k_{1}\cdot k_{3}\xi_{1}\cdot\xi_{3}\xi_{2}\cdot k_{3}+\alpha k_{2}\cdot k_{3}\xi_{2}\cdot\xi_{3}\xi_{1}.k_{2})$ (110)
and
$S^{2}_{0,0}(\bar{z}_{ij})$ $\simeq 0\times q^{0}-8\pi^{2}q+16\pi^{2}q^{2},$ (111)
$S^{2}_{1/2,0}(\bar{z}_{ij})$ $\simeq-\pi^{2}+0\times q+0\times q^{2},$ (112)
$S_{0,0}\left(\bar{z}_{12}\right)S_{0,0}\left(\bar{z}_{31}\right)S_{0,0}\left(\bar{z}_{23}\right)$ $\simeq 0\times q^{0}-8\pi^{3}i\times q^{1}+0\times q^{2}+$... (113)
$S_{1/2,0}\left(\bar{z}_{12}\right)S_{1/2,0}\left(\bar{z}_{31}\right)S_{1/2,0}\left(\bar{z}_{23}\right)$ $\simeq-i\pi^{3}\times q^{0}+0\times q^{1}+0\times q^{2}+$... (114)
Before evaluating the integral, we observe that not all terms in the above equations contribute to the renormalization of the three-gluon tree-level vertex. However, it is nontrivial to determine whether the $l$-dependent terms affect the physical running. Therefore, it is more appropriate to study the typical integrals over the loop momentum and the fundamental region of $\tau$.
We observe that in the non-pinched region,
$|\mathcal{J}_{3}|^{2}$ $=e^{-\alpha^{\prime}\pi\tau_{2}l^{2}-2\pi\alpha^{\prime}\sum_{i}l\cdot k_{i}y_{i}+\pi\alpha^{\prime}\sum_{i<j}k_{i}\cdot k_{j}|y_{ij}|}+\mathcal{O}(q^{2})$
$=e^{-\pi\alpha^{\prime}\tau_{2}p^{2}+\pi\alpha^{\prime}\tau_{2}(\sum_{i}k_{i}\hat{y}_{i})^{2}+\pi\alpha^{\prime}\tau_{2}\sum_{i<j}k_{i}\cdot k_{j}|\hat{y}_{ij}|}+\mathcal{O}(q^{2}),$ (115)
where we define $p=l+\sum_{i}k_{i}y_{i}$ and $\hat{y}_{i}=\frac{y_{i}}{\tau_{2}}$. For later convenience, we will denote $\tilde{k}=\sum_{a}k_{a}y_{a}$, $\Delta=\pi\alpha^{\prime}\tau_{2}(\sum_{i}k_{i}\hat{y}_{i})^{2}+\pi\alpha^{\prime}\sum_{i<j}k_{i}\cdot k_{j}|\hat{y}_{ij}|$ and
$J_{3}=|\mathcal{J}_{3}|=e^{-\frac{\pi}{2}\alpha^{\prime}\tau_{2}p^{2}+\frac{\pi}{2}\alpha^{\prime}\tau_{2}(\sum_{i}k_{i}\hat{y}_{i})^{2}+\frac{\pi}{2}\alpha^{\prime}\tau_{2}\sum_{i<j}k_{i}\cdot k_{j}|\hat{y}_{ij}|}.$ (116)
Recall that
$(\sum_{i}k_{i}\hat{y}_{i})^{2}=2\sum_{i<j}k_{i}.k_{j}y_{i}y_{j}=-\sum_{i<j}k_{i}.k_{j}y_{ij}^{2},$ (117)
so the standard result $\Delta=\pi\alpha^{\prime}k_{i}.k_{j}\sum_{i<j}(|y_{ij}|-y_{ij}^{2})$ in the conventional RNS formalism is obtained.
Then the most typical integral can be expressed as
$I(n,m)=\int_{\mathcal{F}}d^{2}\tau\int d^{4}ll^{n}\tau_{2}^{m}e^{-\pi\alpha\tau_{2}(l^{2}-\Delta)}.$ (118)
If one only interests in the field limit, $I(n,m)$ can be reduced to the world-line limit and regulated by the conventional dimensional regularization,
$I^{1}(n,m)$ $=\int_{0}^{\infty}d\tau_{2}\int d^{4-\epsilon}ll^{n}\tau_{2}^{m}e^{-\pi\alpha\tau_{2}(l^{2}-\Delta)}$
$=\int d^{4-2\epsilon}ll^{n}\frac{m!}{[\pi\alpha^{\prime}(l^{2}-\Delta)]^{m+1}}.$ (119)
Comparing with the Feynman integral in quantum field theory, it is noticeable that $\{\hat{y}_{i}\}$ play the same rules of Feynman parameters.
Besides, we follow the alternative regulation procedure of Green:2008uj and introduce a new cutoff $L$ in the fundamental region as shown in Fig. 2, which is used to extract the field-theory limit as $L\!\rightarrow\!\infty\hskip 0.85358pt$. The physical result should be $L$-independent. Any $L$-dependent term obtained from the point particle region $L,\infty]$ of the integral will be cancelled by another $L$-dependent term from the stringy region (marked by pink color in Fig. 2 of the integral. The $L$-regulated integral is given by
$I^{2}(n,m)$ $=\int_{L}^{\infty}d\tau_{2}\int d^{4}ll^{n}\tau_{2}^{m}e^{-\pi\alpha\tau_{2}(l^{2}-\Delta)}.$ (120)
The more comparison between $L$-regularization and dimensional regularization can be found at Sec.B. Ultimately, these two regularization methods are demonstrated to be equivalent within the theoretical framework of renormalization.
[…]
((a))
Figure 2: The fundamental region is shown as the regions where $|\tau|\hskip-0.85358pt\!>\hskip-0.85358pt\!1$ and $\frac{1}{2}\!\!<\!\tau_{1}\!\!<\!\frac{1}{2}\hskip 0.85358pt$. A cutoff $L\!\!\gg\!\!1$ on the $\tau_{2}$ axis is introduced for the later calculation. The $\tau_{2}\!>\!L$ region corresponds to the region of field-theory limit and $\tau_{2}\hskip-0.85358pt\!<\!L$ is the stringy region.
In addition, it should be stressed that the both $I^{1}(n,m)$ and $I^{2}(n,m)$ converge for $l^{2}-\Lambda>0$; otherwise these integrals diverge. The traditional strategy involves evaluating the integral in the momentum region where convergence is guaranteed and then analytically continuing the result to the divergent region.This procedure is universally valid in the low energy limit $\alpha^{\prime}s_{ij}\rightarrow 0$. For finite $\alpha^{\prime}s_{ij}$, however, contour deformation of the integral becomes necessary. And a rigorous framework is developed in Refs.Witten:2013pra ; Eberhardt:2023xck .
Notice that $m=2$ for the non-pinched diagram and $n\leq 3$ for three-gluon-one-loop amplitude. In the way,
$I(0,2)\sim\frac{1}{\Delta},$ (121)
which leads to IR divergence but UV divergence. And $I(1,2)=I(3,2)=0$ because of the reverse symmetry $l\rightarrow-l$. Only
$I^{1}(2,2)$ $=\frac{2}{(\pi\alpha^{\prime})^{3}\Gamma(3)}i\pi^{2-\epsilon}(2-\epsilon)(\frac{1}{\epsilon}-\gamma)(1-\epsilon\log\Delta)+\mathcal{O}(\epsilon^{1})$ (122)
provides non-zero UV correction. As a consequence, only terms which are proportional to $p^{2}$ in $\Lambda^{1}(\bar{z})$ make non-zero contribution in the non-pinched region, which is given by
$\Lambda^{1}_{\text{red}}(\bar{z})=\alpha^{\prime}(2\pi i)^{3}\frac{1}{4-2\epsilon}l^{2}\xi_{1}.\xi_{2}\xi_{3}.k_{1}(1+y_{12})+\text{cycle}.$ (123)
Here we use the effective replacement $l^{\mu}l^{\nu}\rightarrow\frac{1}{d}l^{2}g^{\mu\nu}$ in the loop momentum integral. Such that
$I_{R}^{\text{scalar}}\sim$ $n_{\text{s}}\Lambda^{1}_{\text{red}}(\bar{z})J_{3},$ (124)
$I_{R}^{\text{vector}}\sim$ $(d-2)\Lambda^{1}_{\text{red}}(\bar{z})J_{3},$ (125)
$I_{R}^{\text{fermion}}\sim$ $-2n_{f}5\Lambda^{1}_{\text{red}}(\bar{z})J_{3}.$ (126)
On the other hand,
$I_{L}^{\text{color}}$ $\sim 4\pi^{3}iC_{\text{SO}(32)}J_{3}.$ (127)
After subtracting the constant and divergent part of $I^{1}(2,2)$, the renormalized non-pinched diagram is given by
$\mathcal{A}^{\text{NP}}_{3(1)}=$ $\frac{~\hskip 0.85358pt\hat{g}^{3}\!}{\,(2\pi)^{3}V\,}2\times\int_{0}^{\hat{y}_{3}}d\hat{y}_{2}\int_{0}^{1}d\hat{y}_{3}\frac{\pi}{8}\xi_{1}\cdot\xi_{2}\xi_{3}\cdot k_{1}(1+y_{12})$
$(d-2+n_{s}-2n_{f})(-\log(\Delta/\mu^{2}))+\text{cycle}.$ (128)
Here factor $2$ accounts the contribution from another alternative order $0=y_{1}<y_{3}<y_{2}$. Evaluating the residue integral over the Feynman parameters $\hat{y}_{2}$ and $\hat{y}_{3}$ is highly non-trivial. Fortunately, since the beta function depends solely on the renormalization scale $\mu$ , we focus on the scale-dependent part of $\mathcal{A}^{\text{NP}}_{3(1)}$ is given by
$\mathcal{A}^{\text{NP}}_{3(1)}(\mu)=\frac{~\hskip 0.85358pt\hat{g}^{3}\!}{\,(2\pi)^{3}V\,}\xi_{1}\cdot\xi_{2}\xi_{3}\cdot k_{1}\frac{\pi}{2\times 3}(d-2+n_{s}-2n_{f})\log\mu+\text{cycle}.$ (129)
This expression reveals the exact cancellation in supersymmetrical theories, where $d-2+n_{s}-2n_{f}=0$. It explains pinched diagrams alone suffice to compute the full beta function of $N=1$ orbifolded heterotic string in Ref.Minahan:1987ha .
3.3.2 Pinched diagram
There are three pinched diagrams or regions,
$\{|z_{12}|\sim\frac{1}{\tau_{2}},|z_{23}|\sim\frac{1}{\tau_{2}},|z_{31}|\sim\frac{1}{\tau_{2}}\}.$ (130)
We focus on the first region $|z_{12}|\sim\frac{1}{\tau_{2}}$ and $|z_{23}|\sim\tau_{2}$ at this subsection. After gauge fixing $z_{1}=0$, $|z_{12}|\sim\frac{1}{\tau_{2}}$ indicates two cases $|z_{2}|\sim\frac{1}{\tau_{2}}$ and $|z_{2}-\tau_{2}|\sim\frac{1}{\tau_{2}}$. The second case arises because of the periodic condition on the torus. When $z_{1}$ and $z_{2}$ pinch, bothe $\partial G(z_{12})$ and $S_{\delta}(z_{12})$ scale as $\frac{1}{z_{12}}$ for any spin structure $\delta$. However, this scaling doesn’t imply this amplitude diverges. We change variables of the moduli integral $d^{2}z_{2}=dx_{2}dy_{2}$ to polar coordinates $\rho d\rho d\theta$, where $\rho=|z_{2}|$ and $\theta\in[0,2\pi]$. Substituting, we obtain
$A_{3(1)}^{\text{P}}\sim\int_{0}^{L}\rho d\rho d\theta\frac{1}{\rho^{2}}\rho^{\alpha^{\prime}k_{1}\cdot k_{2}}=\frac{2\pi L^{\alpha^{\prime}k_{1}\cdot k_{2}}}{\alpha^{\prime}k_{1}\cdot k_{2}}=\frac{2\pi}{\alpha^{\prime}k_{1}\cdot k_{2}}.$ (131)
Here $L$ is a order one free number and in the second equation, we take the limit $\alpha^{\prime}\rightarrow 0$. So as shown as Fig.1, the pinching creates a pole $\frac{2\pi}{\alpha^{\prime}k_{1}\cdot k_{2}}$. So we only need to consider the residue at $z_{ij}\rightarrow 0$.
In the way, it is easy to evaluate the residues at $z_{12}\rightarrow 0$ as
$\text{Res}_{z_{12}}\Phi(\alpha,\beta,z)$ $=S_{\alpha\beta}\hskip-0.85358pt(z_{13})S_{\alpha\beta}\hskip-0.85358pt(z_{31}),$ (132)
$\hskip-25.60747pt\text{Res}_{\bar{z}_{12}}\Lambda^{1}(\bar{z})=$ $\alpha^{\prime}(\xi_{1}\cdot k_{2})(\xi_{2}\cdot k_{3}(\bar{\partial}G_{23})-2\pi il\cdot\xi_{2})\left(-2\pi il\cdot\xi_{3}\right)$
$+\alpha^{\prime}(\xi_{1}\cdot k_{2}(\bar{\partial}G_{31})-2\pi il\cdot\xi_{1})(\xi_{2}\cdot k_{3})\left(-2\pi il\cdot\xi_{3}\right),$ (133)
$\hskip-25.60747pt\text{Res}_{\bar{z}_{12}}\Lambda^{2}(\alpha_{R},\beta_{R},\bar{z})=$ $\alpha^{\prime}\Big[-(k_{1}\cdot k_{3}\xi_{1}\cdot\xi_{3}+\xi_{1}\cdot k_{2}\xi_{3}\cdot k_{1})\xi_{2}\cdot k_{3}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{13}\right)$
$-(k_{2}\cdot k_{3}\xi_{2}\cdot\xi_{3}+\xi_{2}\cdot k_{3}\xi_{3}\cdot k_{1})\xi_{1}\cdot k_{2}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{23}\right)\Big],$ (134)
$\text{Res}_{\bar{z}_{12}}\Lambda^{3}(\alpha_{R},\beta_{R},\bar{z})$ $=\alpha^{\prime}S_{\alpha_{R}\beta_{R}}\left(\bar{z}_{13}\right)S_{\alpha_{R}\beta_{R}}\left(\bar{z}_{23}\right)$
$(+2\alpha^{\prime}\xi_{1}\cdot k_{2}\xi_{2}\cdot k_{3}\xi_{3}\cdot k_{1}+\alpha k_{1}\cdot k_{2}\xi_{1}\cdot\xi_{2}\xi_{3}\cdot k_{1}$
$+\alpha^{\prime}k_{1}\cdot k_{3}\xi_{1}\cdot\xi_{3}\xi_{2}\cdot k_{3}+\alpha k_{2}\cdot k_{3}\xi_{2}\cdot\xi_{3}\xi_{1}\cdot k_{2}).$ (135)
We observe that there are some terms contributing to $\alpha^{\prime}F^{3}$ effective vertex again. After removing these redundant terms, the reduced formulas are given by
$\text{Res}_{\bar{z}_{12}}\Lambda^{1}_{\text{red}}(\bar{z})$
$=$ $\alpha^{\prime}(\xi_{1}\cdot k_{2})(2\pi ip\cdot\xi_{2})\left(2\pi ip\cdot\xi_{3}\right)+\alpha^{\prime}(2\pi ip\cdot\xi_{1})(\xi_{2}\cdot k_{3})\left(2\pi ip\xi_{3}\right),$ (136)
$\text{Res}_{\bar{z}_{12}}\Lambda^{2}_{\text{red}}(\alpha_{R},\beta_{R},\bar{z})+\text{Res}_{\bar{z}_{12}}\Lambda^{3}_{\text{red}}(\alpha_{R},\beta_{R},\bar{z})$
$=$ $\alpha^{\prime}S_{\alpha_{R}\beta_{R}}\left(\bar{z}_{13}\right)S_{\alpha_{R}\beta_{R}}\left(\bar{z}_{23}\right)(\alpha k_{1}\cdot k_{2}\xi_{1}\cdot\xi_{2}\xi_{3}\cdot k_{1}).$ (137)
Besides, once $|z_{12}|\sim\frac{1}{\tau_{2}}$ and $z_{1}=0$, the Koba-Nilsen factor is reduced to be
$\lim_{z_{12}\rightarrow 0}|\mathcal{J}_{3}|^{2}$ $=e^{-\pi\alpha^{\prime}\tau_{2}(l^{2}-\Delta^{\prime}_{12})}+\mathcal{O}(q^{2}),$ (138)
$\Delta^{\prime}_{12}$ $=\lim_{z_{12}\rightarrow 0}\Delta=\pi\alpha^{\prime}k_{1}\cdot k_{2}(y_{3}^{2}-|y_{3}|).$ (139)
So we have
$\text{Res}_{z_{12}}I_{L}^{\text{color}}\sim$ $4\pi^{2}C_{\text{SO}(32)}J_{3}^{(12)},$ (140)
$\text{Res}_{z_{12}}I_{R}^{\text{scalar}}\sim$ $-4\pi^{2}\alpha^{\prime}n_{s}\frac{l^{2}}{4-2\epsilon}(\xi_{2}\cdot\xi_{3}\xi_{1}\cdot k_{2}+\xi_{1}\cdot\xi_{3}\xi_{2}\cdot k_{3})J_{3}^{(12)},$ (141)
$\text{Res}_{z_{12}}I_{R}^{\text{vector}}\sim$ $-4\pi^{2}\alpha^{\prime}(d-2)\frac{l5^{2}}{4-2\epsilon}(\xi_{2}\cdot\xi_{3}\xi_{1}\cdot k_{2}+\xi_{1}\cdot\xi_{3}\xi_{2}\cdot k_{3})J_{3}^{(12)}$
$-\alpha^{\prime}8\pi^{2}(k_{1}\cdot k_{2}\xi_{1}\cdot\xi_{2}\xi_{3}\cdot k_{1})J_{3}^{(12)},$ (142)
$\text{Res}_{z_{12}}I_{R}^{\text{fermion}}\sim$ $8\pi^{2}\alpha^{\prime}n_{f}\frac{l^{2}}{4-2\epsilon}(\xi_{2}\cdot\xi_{3}\xi_{1}\cdot k_{2}+\xi_{1}\cdot\xi_{3}\xi_{2}\cdot k_{3})J_{3}^{(12)}$ (143)
$+\alpha^{\prime 2}2\pi^{2}n_{f}(k_{1}\cdot k_{2}\xi_{1}\cdot\xi_{2}\xi_{3}\cdot k_{1})J_{3}^{(12)},$ (144)
where $J_{3}^{(12)}=\lim_{z_{12}\rightarrow 0}|\mathcal{J}_{3}|$. Hence there are only two integral associating to the pinched diagrams,
$I(2,1)$ $=\frac{-i\Gamma(\epsilon-1)}{\alpha^{\prime 2}\pi^{\epsilon}\Gamma(2)}(2-\epsilon)(\Delta^{\prime}_{12})^{1-\epsilon},$ (145)
$I(0,1)$ $=\frac{i\Gamma(\epsilon)}{\alpha^{\prime 2}\pi^{\epsilon}\Gamma(2)}(\frac{1}{\Delta^{\prime}_{12}})^{\epsilon}.$ (146)
In the way, one can check that
$A_{3(1)}^{\text{P}}(|z_{12}\sim\frac{1}{\tau_{2}}|,\mu)=$ $-\frac{\pi}{2}\frac{~\hskip 0.85358pt\hat{g}^{3}\!}{\,(2\pi)^{3}V\,}\zeta_{1}\cdot\zeta_{2}\zeta_{3}\cdot k_{1}C_{SO(32)}\log(\mu)(4-n_{f}),$
$-\frac{\pi}{2}\frac{~\hskip 0.85358pt\hat{g}^{3}\!}{\,(2\pi)^{3}V\,}\zeta_{2}\cdot\zeta_{3}\zeta_{1}\cdot k_{2}C_{SO(32)}\log(\mu)(\frac{d-2}{12}+\frac{n_{s}}{12}-\frac{2n_{f}}{12}),$
$-\frac{\pi}{2}\frac{~\hskip 0.85358pt\hat{g}^{3}\!}{\,(2\pi)^{3}V\,}\zeta_{3}\cdot\zeta_{1}\zeta_{2}\cdot k_{3}C_{SO(32)}\log(\mu)(\frac{d-2}{12}+\frac{n_{s}}{12}-\frac{2n_{f}}{12}).$ (147)
Sum over the three pinched regions,
$A_{3(1)}^{\text{P}}\sim$ $-\frac{\pi}{2}\frac{~\hskip 0.85358pt\hat{g}^{3}\!}{\,(2\pi)^{3}V\,}C_{SO(32)}(K[1,2;3]+\text{cycle})\log(\mu)(4-n_{f}+\frac{d-2}{6}+\frac{n_{s}}{6}-\frac{2n_{f}}{6}),$ (148)
where $K[1,2;3]\equiv\zeta_{1}\cdot\zeta_{2}\zeta_{3}\cdot k_{1}$.
Accompanying $A_{3(1)}^{\text{NP}}$, it is easy to verify that when $d=4$,
$A_{3(1)}\sim-\frac{i\hat{g}^{3}}{16\pi^{2}V}$ $C_{SO(32)}(K[1,2;3]+\text{cycle})\log(\mu)(\frac{11}{3}-\frac{2}{3}\times n_{f}-\frac{n_{s}}{6}).$ (149)
It corresponds to the beta function given by Eq.(1). So far, we obtain the same results as the quantum field theory with the decomposed string amplitude, which are almost model-independent [^7] The kinetic part is model-independent but the color part is non-trivial to generalized to any Lie group..
3.4 Gravitational beta function
In this section, we shift our focus to the gravitational beta function. Owing to the loop level double copy structure inherent from the chiral-splitting effective formalism, it is sufficient to replace $\Phi_{\alpha_{L}\beta_{L}}$ with $\Lambda^{B}(z)$ derived in Sec.2. It is surprise that we don’t need to care how to split $W(z)=\Lambda^{B}(z)$ into components, the tensor structure of this integral already forces gravitational coupling beta function to vanish.
Due to $l\cdot\zeta_{i}$ factors provided by $\Lambda^{B}_{\text{red}}$, the following effective replacements in the loop momentum integrand should be taken into account,
$l\cdot\xi_{i}l\cdot\zeta_{j}$ $=\frac{l^{2}}{d}\xi_{i}\cdot\zeta_{j},$ (150)
$l\cdot\xi_{i}l\cdot\xi_{j}l\cdot\zeta_{n}l\cdot\zeta_{m}$ $=\frac{l^{4}}{d(d+2)}(\xi_{i}\cdot\xi_{j}\zeta_{n}\cdot\zeta_{m}+\xi_{i}\cdot\zeta_{n}\xi_{j}\cdot\zeta_{m}+\xi_{i}\cdot\zeta_{m}\xi_{j}\cdot\zeta_{n})$ (151)
It is noticeable that there is no $\xi_{i}.\xi_{j}$ in Eq.(3.3.1) as well as $\Lambda^{B}(z)$ in the non-pinched region. Therefore,
$(\Lambda^{1}(\bar{z})\Lambda^{B}(z))_{\text{red}}=$ $[\alpha^{\prime}(2\pi i)^{3}l\cdot\xi_{1}l\cdot\xi_{2}\xi_{3}\cdot k_{1}(1+y_{12})+\text{cycle}]$
$[\alpha^{\prime}(-2\pi i)^{3}l.\zeta_{1}l\cdot\zeta_{2}\zeta_{3}\cdot k_{1}(1+y_{12})+\text{cycle}].$ (152)
Since the polarization tensors of physical graviton are symmetrical and traceless, we will take $\xi_{i}\cdot\zeta_{i}=0$ in Eq.(150) and Eq.(151). Similarly, $\Lambda^{2}(\alpha_{R},\beta_{R},\bar{z})$ and $\Lambda^{3}(\alpha_{R},\beta_{R},\bar{z})$ either provide a $\xi_{i}\cdot\xi_{j}$, $l\cdot\xi_{i}l\cdot\xi_{j}$ or $\xi_{i}\cdot\xi_{j}l\cdot\xi_{l}$,
$(\Lambda^{2}(\alpha_{R},\beta_{R},\bar{z})\Lambda^{B}(z))_{\text{red}}=$ $\Big[-2\pi i(\alpha^{\prime}k_{1}\cdot k_{2})\xi_{1}\cdot\xi_{2}\xi_{3}\cdot k_{1}(1+y_{12})S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{12}\right)$
$-2\pi i(\alpha^{\prime}k_{1}\cdot k_{3})\xi_{1}\cdot\xi_{3}\xi_{2}\cdot k_{3}(1+y_{31})S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{13}\right)$
$-2\pi i(\alpha^{\prime}k_{2}\cdot k_{3})\xi_{2}\cdot\xi_{3}\xi_{1}\cdot k_{2}(1+y_{23})S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{23}\right)\Big]$
$[\alpha^{\prime}(-2\pi i)^{3}l\cdot\zeta_{1}l\cdot\zeta_{2}\zeta_{3}\cdot k_{1}(1+y_{12})+\text{cycle}]$
$+\Big[i2\pi(\alpha^{\prime}k_{1}\cdot k_{2})\xi_{1}\cdot\xi_{2}l\cdot\xi_{3}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{12}\right)$
$+i2\pi(\alpha^{\prime}k_{1}\cdot k_{3})\xi_{1}\cdot\xi_{3}l\cdot\xi_{2}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{13}\right)$
$+i2\pi(\alpha^{\prime}k_{2}\cdot k_{3})\xi_{2}\cdot\xi_{3}l\cdot\xi_{1}S^{2}_{\alpha_{R}\beta_{R}}\left(\bar{z}_{23}\right)\Big]$
$\Big[-(-2\pi i)^{3}l\cdot\zeta_{1}\zeta_{2}\cdot k_{3}(1+y_{31})\zeta_{3}\cdot k_{1}(1+y_{12})+\text{cycle}\Big],$ (153)
$(\Lambda^{2}(\alpha_{R},\beta_{R},\bar{z})\Lambda^{B}(z))_{\text{red}}=$ $S_{\alpha_{R}\beta_{R}}\left(\bar{z}_{12}\right)S_{\alpha_{R}\beta_{R}}\left(\bar{z}_{13}\right)S_{\alpha_{R}\beta_{R}}\left(\bar{z}_{23}\right)$
$(\alpha k_{1}\cdot k_{2}\xi_{1}\cdot\xi_{2}\xi_{3}\cdot k_{1}+\alpha^{\prime}k_{1}\cdot k_{3}\xi_{1}\cdot\xi_{3}\xi_{2}\cdot k_{3}+\alpha k_{2}\cdot k_{3}\xi_{2}\cdot\xi_{3}\xi_{1}\cdot k_{2})$
$[\alpha^{\prime}(-2\pi i)^{3}l\cdot\zeta_{1}l\cdot\zeta_{2}\zeta_{3}\cdot k_{1}(1+y_{12})+\text{cycle}].$ (154)
If these integrals doesn’t vanish due to supersymmetry, there are only two kinds of integrals required, which are given by
$I^{1}(4,2)$ $=\frac{-i2\pi^{2-\epsilon}}{\pi^{3}\alpha^{\prime 3}}(2-\epsilon)(3-\epsilon)\frac{\Gamma(\epsilon-1)}{2}\Delta^{1-\epsilon},$ (155)
and $I^{1}(2,2)$ as shown as Eq.(122). It is noticeable that the loop corrections in non-pinched diagram are universally proportional to $s_{ij}\log(C\Delta)$ before the integral over Feynman parameters, where $C$ is a constant dependent on the regularization scheme. The most general renormalization condition can chosen to be
$s_{12}=a_{12}\mu^{2},s_{23}=a_{23}\mu^{2},s_{31}=a_{31}\mu^{2}$ (156)
with $a_{12}+a_{23}+a_{31}=0$ and $a_{ij}\in R$. However, the subtraction procedure remains ambiguous. Conventionally, the renormalized one loop amplitude could be proportional to
$s_{nm}\log(\Delta)-xs_{nm}\log(\sum_{i<j}\pi\alpha^{\prime}(y^{2}_{ij}-|y_{ij}|)a_{ij}\mu^{2})-ya_{nm}\mu^{2}\log(\sum_{i<j}\pi\alpha^{\prime}(y^{2}_{ij}-|y_{ij}|)a_{ij}\mu^{2}),$ (157)
for any $x+y=1$ and $x,y\in R$. Physically, $a_{nm}\mu^{2}\log(\sum_{i<j}\pi\alpha^{\prime}(y^{2}_{ij}-|y_{ij}|)a_{ij}\mu^{2})$ corresponds to the power-law running of gravitational coupling constant while $s_{nm}\log(\sum_{i<j}\pi\alpha^{\prime}(y^{2}_{ij}-|y_{ij}|)a_{ij}\mu^{2})$ generates a loop correction to a higher dimensional operator.
However, the term $a_{nm}\mu^{2}\log(\sum_{i<j}\pi\alpha^{\prime}(y^{2}_{ij}-|y_{ij}|)a_{ij}\mu^{2})$ introduce a dependence of the beta function on the physically in-relevant parameters $a_{ij}$. So the only self-consistent choice is $x=1$ and $y=0$, which ensures the running of physical coupling constant independent on the artificial renormalization conditions. Beyond the consistent arguments of renormalization, this result is also corroborated by the $L$-regularization
$I^{2}(4,2)=\frac{\Gamma(5)}{(\pi\alpha^{\prime})^{4}\Gamma(2)}(\pi\Delta)\int d\alpha^{\prime}(-\log L-\gamma-\log(\pi\alpha^{\prime}\Delta)+\mathcal{O}(L^{1})).$ (158)
Since the string-derived counter-terms are one to one corresponding to the $L$-dependent terms given by $L$-regulation, Eq.(158) demonstrates that all the counter terms provided from string theory should be proportional to $s_{ij}$, which only corresponds to $s_{nm}\log(\sum_{i<j}\pi\alpha^{\prime}\\ (y^{2}_{ij}-|y_{ij}|)a_{ij}\mu^{2})$. Consequently, the non-pinched diagram contributes loop corrections to higher-dimensional operators but does not renormalize the gravitational coupling constant.
Nextly, we consider the $|z_{12}|\sim\frac{1}{\tau_{2}}$ without loss of generality. The residue of $\Lambda^{B}_{\text{red}}(z)$ is given by
$\text{Res}_{z_{12}}\Lambda^{B}_{\text{red}}(z)=\alpha^{\prime}(\xi_{1}\cdot k_{2})(+2\pi il\cdot\xi_{2})\left(+2\pi il\cdot\xi_{3}\right).$ (159)
With Eq.(3.3.2) and Eq.(3.3.2), the relevant integrals to be considered are provided by
$I(4,1)$ $=\frac{i2\pi^{2-\epsilon}}{\pi^{2}\alpha^{\prime 2}}(2-\epsilon)(3-\epsilon)\frac{\Gamma(\epsilon-2)}{1}\Delta^{{}^{\prime}2-\epsilon}_{ij}.$ (160)
and $I(2,1)$ as shown as Eq.(145).
Even though a pole $\frac{2\pi}{\alpha^{\prime}k_{1}\cdot k_{2}}$ arises from pinching $z_{1}$ and $z_{2}$, it is always canceled by the $k_{1}\cdot k_{2}$ induced by OPE or $k_{1}\cdot k_{2}$ contained in $\Delta^{\prime}_{12}$. Finally, the loop correction from the pinched diagram $|z_{12}\sim\frac{1}{\tau_{2}}|$ is proportional to
$M_{3(1)}^{\text{P}}(|z_{12}|\sim\frac{1}{\tau_{2}})\sim k_{1}\cdot k_{2}\log(\alpha^{\prime}k_{1}\cdot k_{2}).$ (161)
The same arguments as those for non-pinched diagram shows that this is not a renormalization of gravitational constant but rather a higher dimensional operator. Similar results should verify for the other pinched diagrams.
In conclusion, perturbative gravitational renormalization is absent in string models compactified from the heterotic string,
$\beta_{\text{gravity}}=0.$ (162)
However, it remains unclear whether quantum field effects can modify the gravitational coupling through quantum corrections to the dilaton vacuum once more realistic moduli stabilization Grana:2005jc ; Douglas:2006es is taken into account. Such effects cannot be computed within perturbative string theory, highlighting the need for non-perturbative approaches.
3.5 Remark on the gravitational correction to gauge beta function
Notice that in Sec.3.3, we didn’t consider $I_{L}^{\text{vector}}$ and $I_{L}^{\text{scalar}}$. According to Table.2, these two term reflect the correction from 10d gravity sector including 4d gravity sector and other extra-dimensional components, for example, gravi-photon. Because of $W_{0}=0$,
$I_{L}^{\text{vector}}$ $=\sum_{\mu=1}^{d}T_{\mu}^{(2)}|_{\text{multi-linear}},$ (163)
$I_{L}^{\text{scalar}}$ $=\sum_{\mu=d}^{n_{s}+d}T_{\mu}^{(2)}|_{\text{multi-linear}}.$ (164)
While the external states are 4d massless gluon without extra-dimensional momentum and polarization vectors, it is straight-forward to obtain
$I_{L}^{\text{scalar}}=0.$ (165)
Thus we reach the first conclusion that there is no correction to gauge beta function from other extra-dimensional components of higher dimensional gravity.
Then Eq.101,Eq.102 and Eq.103 shows that $\partial G(z)$ and $\partial^{2}G(z)$ have no contribution except $G(z)$. Especially, non-trivial part of $I_{L}^{\text{vector}}$ comes from the $q$-expansion of $\mathcal{J}_{3}$. One instantly realize that
  1. 1.
    For non-pinched diagram, the momentum conservation and on-shell condition cause
    $I_{L}^{\text{vector}}\sim\sum_{i<j}k_{i}\cdot k_{j}=0.$ (166)
  2. 2.
    For pinched diagram where $z_{ij}\rightarrow 0$,
    $\text{Res}_{z_{ij}\rightarrow 0}I_{L}^{\text{vector}}\sim-k_{i}\cdot k_{j}J_{3}^{(ij)}.$ (167)
The gravitational sector, including the graviton, antisymmetric tensor, and dilaton, does indeed receive non-zero quantum corrections to the three-gluon one-loop diagram. However, as argued in Section 3.4 regarding the renormalization conditions for the gravitational beta function, this loop correction generates higher-dimensional operators rather than renormalizing the gauge coupling. Such that
$\Delta_{\text{gravity}}\beta_{\text{gauge}}=0.$ (168)
Thanks to the internal-line double copy framework, our analysis and calculations verify these results across a broad class of theories derived from heterotic string theory, in contrast to case-by-case Feynman diagram computations. This result also relies on the perturbative approach and the non-perturbative correction from gravity to gauge coupling remains unknown.
4 Summary
In this paper, the double-copy-like decomposition for the internal line, as introduced in Refs.Tourkine:2012vx ; Geyer:2015jch , is developed and generalized to bosonic (Sec.2.1) and heterotic (Sec.2.2) string theories in Sec.2. Compared to the formulation in Refs.Tourkine:2012vx ; Geyer:2015jch , we employ the chiral-splitting effective formalism to make the double copy structure manifest, and include the chiral Koba-Nielsen factors $\mathcal{J}_{n}$ and $\bar{\mathcal{J}}_{n}$ in the $q\bar{q}$-expansion. These factors are not essential for superstring theories or the supersymmetric right movers of the heterotic string, since $\mathcal{J}_{n}$ and $\bar{\mathcal{J}}_{n}$ contribute non-trivially only for $q^{n\geq 2}$ and $\bar{q}^{n\geq 2}$, whose partition functions exhibit poles of the form $\sim\{1/q,1/\bar{q}\}$. However, the bosonic string partition function contains poles $\sim\{1/q^{2},1/\bar{q}^{2}\}$, and the left movers of the heterotic string feature poles $\sim\{1/q^{2},1/\bar{q}^{2},1/q,1/\bar{q}\}$, making the chiral Koba-Nielsen factors non-negligible. Furthermore, the fermionic realization of the color degrees of freedom in the heterotic string is emphasized, and the corresponding decomposition is formulated in Sec.2.2.
In Sec.3, we applied the referenced internal line double copy to compute the physical beta functions for the gauge coupling constant and the gravitational coupling constant. The detailed scattering integrands were presented in Sec.3.1. Thanks to the chiral-splitting effective formalism, the double copy structure of the integrands became manifest. One can observe that the difference between these two integrands lies in their left-moving parts, which are $\Phi_{\alpha_{L}\beta_{L}}$ and $\Lambda^{b}$, respectively. In the quantum field theory computation of the beta functions, multiple Feynman diagrams contribute. These diagrams are realized in the string framework through different regions of moduli space, and the corresponding decomposition was discussed in Sec.3.2.
Later, the most general QCD beta function was reproduced in Sec.3.3. Unlike previous approaches that extract gauge beta functions from string amplitudes using specific models, our method does not rely on any specialized model and directly yields the general result. In addition, we explored the possible gravitational beta function in Sec.3.4. It was found that $\Lambda^{b}$ contains a more intricate dependence on the loop momentum $l$, leading to nontrivial Mandelstam factors $s_{ij}$ in the final expression. A careful analysis of renormalization revealed that these quantum corrections involving additional Mandelstam factors should be interpreted as higher-dimensional operators rather than as renormalizations of the original gravitational coupling constant.
Finally, we studied the gravitational correction to the three-gluon one-loop amplitude in Sec.3.5. This correction arises solely from the chiral Koba-Nielsen factors $\mathcal{J}_{3}$ and is always accompanied by Mandelstam variables. As a result, it cancels due to momentum conservation in non-pinched diagrams, but remains non-zero in pinched diagrams. However, a similar renormalization analysis indicates that this contribution should also be interpreted as a higher-dimensional operator, not as a renormalization of the gauge coupling constant.
Acknowledgements.
I sincerely thanks Hong-Jian He for discussing the gravitational correction to gauge coupling constant.
Appendix A Spin Structure Summation
There is $2^{2g}$ different spin structure for $g$ genus Riemann surface. And the final physics result is not dependent on spin structure. Thus it is usually taken to be a drawback of the RNS formalism. GS and pure spinor formalism can yield the same result without spin structure summation.
We recall that the world sheet fermion two point correlation function for given spin structure $v$ is given by
$\langle\psi^{M}(z)\psi^{N}(0)\rangle_{v}\sim\eta^{MN}\partial G(z),v=1;\eta^{MN}S_{v}(z),v=2,3,4.$ (169)
Here we define
$S_{v}(z)$ $=\frac{\theta_{1}^{\prime}(0)\theta_{v}(z)}{\theta_{v}(0)\theta_{1}(z)},\nu=2,3,4$ (170a)
$G(z,\bar{z})$ $=-\frac{1}{2}\ln|\frac{\theta_{1}(z,\tau)}{\theta^{\prime}_{1}(0,\tau)}|^{2}+\frac{(\Im mz)^{2}}{4\pi\tau_{2}}$ (170b)
Then we consider the following Kronecker–Eisenstein series and the associated double period function,
$F(z,\eta,\tau)$ $=\frac{\theta_{1}^{\prime}(0)\theta_{1}(z+\eta)}{\theta_{1}(\eta)\theta_{1}(z)},$ (171)
$\Omega(z,\eta,\tau)$ $=e^{2\pi i\eta\frac{Imz}{\tau_{2}}}F(z,\eta,\tau).$ (172)
It is easy to notice that
$F(z_{1}-z_{2},\eta,\tau)F(z_{2}-z_{3},\eta,\tau)...F(z_{n}-z_{1},\eta,\tau)$
$=$ $\Omega(z_{1}-z_{2},\eta,\tau)\Omega(z_{2}-z_{3},\eta,\tau)...\Omega(z_{n}-z_{1},\eta,\tau)$
$=$ $\sum_{n}\eta^{n}V_{n}(z_{1},...,z_{n},\tau).$ (173)
Here $V_{n}(z_{1},...,z_{n},\tau)$ is a function of $f_{i}(z,\tau)$.
Notice that theta functions can be mapped to each other with the following relations,
$\theta_{2}(z+\frac{1}{2},\tau)$ $=-\theta_{1}(z,\tau),$ (174)
$\theta_{4}(z+\frac{\tau}{2},\tau)$ $=ie^{-i\pi z}q^{-\frac{1}{8}}\theta_{1}(z,\tau),$ (175)
$\theta_{4}(z+\frac{1}{2},\tau)$ $=\theta_{3}(z,\tau),$ (176)
$\theta_{3}(z+\frac{\tau}{2},\tau)$ $=e^{-i\pi z}q^{-\frac{1}{8}}\theta_{2}(z,\tau).$ (177)
So that we define
$w_{2}$ $=\frac{1}{2},$ (178)
$w_{3}$ $=-\frac{1+\tau}{2},$ (179)
$w_{4}$ $=\frac{\tau}{2}$ (180)
and
$\Omega(\vec{z}_{n},w_{v},\tau)=$ $\sum_{v}S_{v}(z_{1}-z_{2},\tau)S_{v}(z_{2}-z_{3},\tau)...S_{v}(z_{n}-z_{1},\tau)$
$=$ $\sum_{v}F(z_{1}-z_{2},w_{v},\tau)F(z_{2}-z_{3},w_{v},\tau)...F(z_{n}-z_{1},w_{v},\tau)$
$=$ $\sum_{v}\Omega(z_{1}-z_{2},w_{v},\tau)\Omega(z_{2}-z_{3},w_{v},\tau)...\Omega(z_{n}-z_{1},w_{v},\tau).$ (181)
Finally, we need to evaluate
$\mathcal{G}_{n}(z_{1},...,z_{n},\tau)=\sum_{v}(-1)^{v+1}(\frac{\theta_{v}(0,\tau)}{\theta^{\prime}_{1}(0,\tau)})^{4}S_{v}(z_{1}-z_{2},\tau)S_{v}(z_{2}-z_{3},\tau)...S_{v}(z_{n}-z_{1},\tau).$ (182)
Using
$(-1)^{v+1}(\frac{\theta_{v}(0,\tau)}{\theta^{\prime}_{1}(0,\tau)})^{4}=$ $\frac{1}{(e_{1}-e_{2})(e_{1}-e_{3})},v=2$ (183)
$(-1)^{v+1}(\frac{\theta_{v}(0,\tau)}{\theta^{\prime}_{1}(0,\tau)})^{4}=$ $\frac{1}{(e_{2}-e_{1})(e_{2}-e_{3})},v=3$ (184)
$(-1)^{v+1}(\frac{\theta_{v}(0,\tau)}{\theta^{\prime}_{1}(0,\tau)})^{4}=$ $\frac{1}{(e_{3}-e_{1})(e_{3}-e_{2})},v=4$ (185)
where $e_{v}=\wp(w_{v})$. And
$\Omega(\vec{z}_{n},\eta,\tau)=\sum_{k}^{n-2}\frac{(-1)^{n-k}}{(n-k-1)!}(\wp^{(n-k-2)}(\eta)-\hat{G}_{n-k-2})V_{k}+V_{n}.$ (186)
In addition, $\partial\wp(z)$ vanishes on $z=w_{v}$ and the differential equation So $\mathcal{G}_{n}$ can be expressed as $V_{k}$ and $e_{v}$ in general. Some useful results in literature are
$\mathcal{G}_{n\leqslant 3}=0\,$, (187a)
$\mathcal{G}_{4\leqslant n\leqslant 7}=V_{n}\,.$ (187b)
Appendix B $L$-regularization
We consider the following integrals
$I_{1}(2n,m)$ $=\int_{0}^{\infty}d\tau_{2}\int d^{4-\epsilon}ll^{2n}\tau_{2}^{m}e^{-\pi\alpha\tau_{2}(l^{2}-\Delta)},$ (188)
$I_{2}(2n,m)$ $=\int_{L}^{\infty}d\tau_{2}\int d^{4}ll^{2n}\tau_{2}^{m}e^{-\pi\alpha\tau_{2}(l^{2}-\Delta)}.$ (189)
One can first consider the integral over loop momemtum. If we set $n=0$ for simplicity,
$\int d^{d}le^{-\pi\alpha^{\prime}\tau_{2}l^{2}}=(\pi\alpha^{\prime}\tau_{2})^{-\frac{d}{2}}.$ (190)
And for $n\neq 0$, one can find the following recursive realtion,
$\int d^{d}ll^{2n}e^{-\pi\alpha^{\prime}\tau_{2}l^{2}}=$ $\int d^{d}l(-\frac{1}{\pi\alpha^{\prime}}\frac{\partial}{\partial\tau_{2}})^{n}e^{-\pi\alpha^{\prime}\tau_{2}l^{2}}$
$=$ $(-1)^{n}\frac{\Gamma(\frac{d}{2}+n+1)}{\Gamma(\frac{d}{2})}(\pi\alpha^{\prime}\tau_{2})^{-\frac{d}{2}-n}.$ (191)
So that to prove $I_{1}(n,m)$ and $I_{2}(n,m)$ can be rewritten as
$I^{1}(2n,m)$ $=(-1)^{n}\frac{\Gamma(n+3-\epsilon)}{(\pi\alpha^{\prime})^{n+2-\epsilon}\Gamma(2-\epsilon)}\int_{0}^{\infty}d\tau_{2}\tau_{2}^{m-n-2+\epsilon}e^{\pi\alpha\tau_{2}\Delta},$ (192)
$I^{2}(2n,m)$ $=(-1)^{n}\frac{\Gamma(n+3)}{(\pi\alpha^{\prime})^{n+2}\Gamma(2)}\int_{L}^{\infty}d\tau_{2}\tau_{2}^{m-n-2}e^{\pi\alpha\tau_{2}\Delta},$ (193)
where we assume $\Delta<0$. The directly calculation by Mathematica shows that
$\int_{0}^{\infty}d\tau_{2}\tau_{2}^{-n^{\prime}+\epsilon}e^{\pi\alpha\tau_{2}\Delta}$ $=(\pi\Delta)^{n^{\prime}-1}(\int d\alpha^{\prime})^{n^{\prime}-1}\Gamma(\epsilon)(\pi\alpha^{\prime}\Delta)^{-\epsilon}$
$=(\pi\Delta)^{n^{\prime}-1}(\int d\alpha^{\prime})^{n^{\prime}-1}(\frac{1}{\epsilon}-\gamma-\log{\pi\alpha^{\prime}\Delta}+\mathcal{O}(e)),$ (194)
$\int_{L}^{\infty}d\tau_{2}\tau_{2}^{-n^{\prime}}e^{\pi\alpha\tau_{2}\Delta}$ $=(\pi\Delta)^{n^{\prime}-1}(\int d\alpha^{\prime})^{n^{\prime}-1}(-\log(L)-\gamma-\log{\pi\alpha^{\prime}\Delta}+\mathcal{O}(L)).\$(195)
We see the formal mappling between $\frac{1}{\epsilon}\leftrightarrow-\log(L)$ and $\epsilon\leftrightarrow L$. However, because there are also $\epsilon$ dependence on $\frac{\Gamma(n+3-\epsilon)}{(\pi\alpha^{\prime})^{n+2-\epsilon}\Gamma(2-\epsilon)}$ in Eq.(192), we have
$I^{1}(2n,m)-I^{2}(2n,m)=(\pi\alpha^{\prime}\Delta)^{m-n-3}\times\text{Const}.$ (196)
While any constant can be absorbed into the Logritham function with
$\log(\pi\alpha^{\prime}\Delta)+\log(C)=\log(C\pi\alpha^{\prime}\Delta).$ (197)
In this sense, $I^{1}(2n,m)$ and $I^{2}(2n,m)$ are equivalent to each other as regularization approaches but in mathematical senses.

### v2

@@
-Referenced internal-line double copy and application to gauge and gravitational beta functions
+Referenced internal-line decomposition of one-loop string integrand and application to gauge and gravitational beta functions
 Abstract
-Following the approach of Refs.
-Tourkine:2012vx ; Geyer:2015jch , the double-copy-like decomposition of exchanged internal states in the world-line limit of one-loop string amplitudes is systematically formulated and generalized to both bosonic and heterotic string theories.
+Following the idea of Refs.
+Tourkine:2012vx ; Geyer:2015jch , the double-copy-like decomposition of one loop string integrand according to the loop-exchanged internal states is systematically reformulated and generalized to both bosonic and heterotic string theories.
 As an application, the one-loop beta functions for the gauge and gravitational coupling constants are investigated by analyzing the low-energy field-theory limit of the corresponding three-point one-loop amplitudes in heterotic string theory under a naive $T^{6}$ compactification.
 Due to supersymmetry, these beta functions vanish trivially.
-However, by decomposing the scattering integrand according to the different internal loop-exchanged states, the most general model-independent results are obtained.
+However, by decomposing the scattering integrand according to the different internal loop-exchanged states, the model-independent results of these beta functions are obtained.
 1 Introduction
-The double copy relation reveals a profound connection between the perturbative scattering amplitudes of gauge theories and gravitational theories, expressed as $\text{GR}=\text{YM}^{2}$ Bern:2019prr .
-Although the tree-level double copy originates from the KLT relations Kawai:1985xq in string theory, recent advances in quantum field theory—such as the 5-loop calculation Bern:2018jmv and the classical solution double copy Monteiro:2014cda —have extended beyond the current reach of perturbative string theory.
+The double copy relation reveals a profound connection between the perturbative scattering amplitudes of gauge theories and gravitational theories, expressed as $\text{GR}=\text{YM}^{2}$.
+The most typical $L$-loop $n$-point scattering amplitude with manifest double copy structure can be written as
+$A^{n}_{L}\sim\int d^{LD}l\sum_{i}\frac{1}{S_{i}}\frac{N_{i}\tilde{N}_{i}}{D_{i}},$ (1)
+where $l$ denotes the loop momentum, $i$ ranges among all the possible $\phi^{3}$ Feynman diagrams, $D_{i}$ denotes the products of corresponding scalar propagators, $S_{i}$ denotes possible symmetrical factors, $N_{i}$ and $\tilde{N}_{i}$ are called BCJ numerators which satisfy the color-kinematic dualityBern:2008qj ; Bern:2019prr .
+Historically, the tree-level double copy originated from the KLT relations Kawai:1985xq in tree-level string amplitude.
+While the celebrated CHY formalismCachazo:2013gna ; Cachazo:2013hca ; Cachazo:2013iea extended the tree level double copy relation to other theories.
+And later ambitwistor stringMason:2013sva ; Geyer:2014fka ; Geyer:2015jch ; Geyer:2019hnn further spelled the loop amplitude as well as the loop level double copy.
+Besides, it is also possible to explore the loop level scattering integrand by the forward limit of tree level scattering integrandHe:2015yua ; Cachazo:2015aol ; Cao:2025ygu .
+By the way, one can also study the field theory double copy by low energy limit of string amplitudeHe:2017spx ; Mafra:2017ioj ; Mizera:2019blq ; Casali:2020knc ; Britto:2021prf ; Balli:2024wje .
+Although recent advances in quantum field theory, such as the 5-loop calculation Bern:2018jmv and the classical solution double copy Monteiro:2014cda , have extended beyond the current reach of perturbative string theory.
 Nevertheless, string amplitudes continue to serve as an important laboratory for studying scattering amplitudes.
-The main focus of research on the perturbative double copy has traditionally been on the correspondence between external states.
+It is noticeable that the information of the loop-exchanged internal states are not manifest in Eq.1.
+In contraction, the BCJ numerators $N_{i}$ and $\tilde{N}_{i}$ are functions of external states.
+So we argue that the main focus of research on the perturbative double copy has traditionally been on the correspondence between external onshell states.
 However, studies about low energy limit of one loop type II string amplitude in Ref.Tourkine:2012vx and similar work on ambitwistor string Geyer:2015jch have opened new avenues for the double copy structure of internal states exchanged at one loop.
@@
 The non-Abelian nature of QCD gauge theory results in a negative beta function, leading to the celebrated asymptotic freedom of the running gauge coupling constant.
-This serves as an ideal test case to examine whether the double copy of the reference internal line yields a correct result.
-Moreover, given the non-Abelian nature of general relativity (GR), it is intriguing to investigate the loop corrections to the gravitational beta function arising from the graviton (spin-2), dilaton (spin-0, scalar), antisymmetric tensor field (spin-0, axial scalar in 4D), and gravitino (spin-3/2).
+This serves as an ideal test case to examine whether the decomposition of the reference internal line yields a correct result.
+Several works in the literature Minahan:1987ha ; DiVecchia:1996uq ; DiVecchia:1996iz derived the gauge beta function from the low-energy limit of string amplitudes.
+However, these calculations are restricted to specialized string models with fixed field content $\{n_{v},n_{s},n_{f}\}$ .
+In contrast, the referenced internal-line decomposition delivers a significant advantage: it directly yields a universal result valid for arbitrary $\{n_{v},n_{s},n_{f}\}$.
+Moreover, given the non-Abelian nature of general relativity (GR), it is intriguing to investigate the loop corrections to the gravitational beta function arising from the graviton (spin-2), dilaton (spin-0, scalar), antisymmetric tensor field (spin-0, axial scalar in 4D) and gravitino (spin-3/2).
+However, due to the complicated Feynman rules, it is very hard to perform such loop level calculation in quantum field theory.
+And the loop-exchanged states information is not clear in the traditional loop-level double copy as shown as Eq.1.
+Fortunately, we shall see that these loop states present as a double copy after using chiral-splitting effective formalism and internal-line decomposition.
 At the same time, one may expect that gravity also induces corrections to the gauge beta function, proportional to the gravitational constant $\kappa$, where $\kappa^{2}=32\pi G$ and $G$ is the Newtonian gravitational constant.
@@
 (3) was first proposed by Robinson and Wilczek Robinson:2005fj , and had resulted in a long time debate Pietrykowski:2006xy ; Toms:2007sk ; Ebert:2007gf ; He:2010mt ; Anber:2010uj ; Ellis:2010rw .
-This issue is also addressed in the present work with referenced internal line double copy.
-Several works in the literature Minahan:1987ha ; DiVecchia:1996uq ; DiVecchia:1996iz derived the gauge beta function from the low-energy limit of string amplitudes.
-However, these calculations are restricted to specialized string models with fixed field content $\{n_{v},n_{s},n_{f}\}$ .
-In contrast, the referenced internal-line double-copy approach delivers a significant advantage: it directly yields a universal result valid for arbitrary $\{n_{v},n_{s},n_{f}\}$.
+This issue is also addressed in the this work with referenced internal line decomposition and double copy.
 This paper is organized as follow.
@@
 One can see that the double copy structure is restored after introducing the loop momentum $l$ with the cost of non-manifest modular symmetry.
-More serious loop level KLT relation Stieberger:2023nol was also formulated under this formalism.
+The loop level KLT relation Stieberger:2023nol was also formulated under the chiral splitting effective formalism.
 Next, we turn to the world-line limit $\tau_{2}\rightarrow\infty$.
@@
 $P_{L}(1/2,1/2,\tau)=1,$ (71)
-which reflects the fact that only massless fermions are exchanged in these diagrams as well as quantum field theory.
+which reflects the fact that only massless fermions are exchanged in these diagrams as well as quantum field theory. The internal-line decomposition is trivial in odd spin structure.
 At the end of this section, we summary the double copy relation of referenced internal-line of heterotic string in Table.2.
@@
 $-2\zeta_{2}\cdot\zeta_{3}\partial^{2}G_{23}\left(\zeta_{1}\cdot k_{2}\left(\partial G_{12}-\partial G_{13}\right)+2\pi il\cdot\zeta_{1}\right).$ (90)
-It is obvious to see that $\Lambda^{b}(z)$ is the complex conjugate of $\Lambda^{1}(\bar{z})$.
+It is obvious to see that $\Lambda^{b}(z)$ is the complex conjugate of $\Lambda^{1}(\bar{z})$ with replacing $\xi_{i}$ by $\zeta_{i}$.
 3.2 Decomposition of moduli space
@@
 $S_{1/2,0}\left(\bar{z}_{12}\right)S_{1/2,0}\left(\bar{z}_{31}\right)S_{1/2,0}\left(\bar{z}_{23}\right)$ $\simeq-i\pi^{3}\times q^{0}+0\times q^{1}+0\times q^{2}+$... (115)
-Before evaluating the integral, we observe that not all terms in the above equations contribute to the renormalization of the three-gluon tree-level vertex.
+Before evaluating this integral, we observe that not all terms in the above equations contribute to the renormalization of the three-gluon tree-level vertex.
 However, it is nontrivial to determine whether the $l$-dependent terms affect the physical running.
@@
 Similar results should verify for the other pinched diagrams.
-In conclusion, perturbative gravitational renormalization is absent in string models compactified from the heterotic string,
+In conclusion, perturbative gravitational renormalization is absent in any string model compactified from the heterotic string,
 $\beta_{\text{gravity}}=0.$ (163)
@@
 Such effects cannot be computed within perturbative string theory, highlighting the need for non-perturbative approaches.
-3.5 Remark on the gravitational correction to gauge beta function
+3.5 The gravitational correction to gauge beta function
 Notice that in Sec.3.3, we didn’t consider $I_{L}^{\text{vector}}$ and $I_{L}^{\text{scalar}}$.
@@
 Then Eq.102,Eq.103 and Eq.104 shows that $\partial G(z)$ and $\partial^{2}G(z)$ have no contribution except $G(z)$.
-Especially, non-trivial part of $I_{L}^{\text{vector}}$ comes from the $q$-expansion of $\mathcal{J}_{3}$.
+Hence the non-trivial part of $I_{L}^{\text{vector}}$ only comes from the $q$-expansion of $\mathcal{J}_{3}$.
 One instantly realize that
   1. 1.
-    For non-pinched diagram, the momentum conservation and on-shell condition cause
+    For non-pinched diagram, the momentum conservation and on-shell conditions lead to
     $I_{L}^{\text{vector}}\sim\sum_{i<j}k_{i}\cdot k_{j}=0.$ (167)
@@
     $\text{Res}_{z_{ij}\rightarrow 0}I_{L}^{\text{vector}}\sim-k_{i}\cdot k_{j}J_{3}^{(ij)}.$ (168)
-The gravitational sector, including the graviton, antisymmetric tensor, and dilaton, does indeed receive non-zero quantum corrections to the three-gluon one-loop diagram.
+It reals that the gravitational sector, including the graviton, antisymmetric tensor, and dilaton, does indeed provide non-zero quantum corrections to the three-gluon one-loop diagram if some supersymmetries are broken.
 However, as argued in Section 3.4 regarding the renormalization conditions for the gravitational beta function, this loop correction generates higher-dimensional operators rather than renormalizing the gauge coupling.
-Such that
+Such that there always is
 $\Delta_{\text{gravity}}\beta_{\text{gauge}}=0.$ (169)
-Thanks to the internal-line double copy framework, our analysis and calculations verify these results across a broad class of theories derived from heterotic string theory, in contrast to case-by-case Feynman diagram computations.
+Thanks to the internal-line decomposition and double copy framework, our analysis and calculations verify these results across a broad class of theories derived from heterotic string theory, in contrast to case-by-case Feynman diagram computations.
 This result also relies on the perturbative approach and the non-perturbative correction from gravity to gauge coupling remains unknown.
@@
 Furthermore, the fermionic realization of the color degrees of freedom in the heterotic string is emphasized, and the corresponding decomposition is formulated in Sec.2.2.
-In Sec.3, we applied the referenced internal line double copy to compute the physical beta functions for the gauge coupling constant and the gravitational coupling constant.
+In Sec.3, we applied the referenced internal line decomposition and double copy to compute the physical beta functions for the gauge coupling constant and the gravitational coupling constant.
 The detailed scattering integrands were presented in Sec.3.1.
@@
 Acknowledgements.
-I sincerely thanks Hong-Jian He for discussing the gravitational correction to gauge coupling constant.
+I sincerely thank Hong-Jian He for suggesting and discussing the problem about loop correction to gauge and gravitational beta functions.
+And the two figures in this paper are recycled from the unpublished draft with Hong-Jian He on string loop correction to gauge beta functions by traditional RNS formalism without internal-line decomposition and chiral splitting effective formalism.
 Appendix A Spin Structure Summation